### v3

# Quasi-linear equation $\Delta_{p}v+av^{q}=0$ on manifolds with integral bounded Ricci curvature and geometric applications

Abstract.
We study nonexistence results and gradient estimates for solutions of
$\Delta_{p}v+av^{q}=0$
defined on complete Riemannian manifolds satisfying a $\chi$-type Sobolev inequality. We establish a Liouville theorem under the assumptions that the underlying manifold $(M,g)$ supports a $\chi$-type Sobolev inequality and that the $L^{\frac{\chi}{\chi-1}}$-norm of $\mathrm{Ric}_{-}(x)$ is bounded above by a constant depending only on $\dim(M)$, the Sobolev constant $\mathbb{S}_{\chi}(M)$, and the volume growth rate of geodesic balls $B_{r}\subset M$. This extends and improves several recent results of Ciraolo, Farina, and Polvara [18]; our approach, however, differs from their “$P$-function” method. In addition, for manifolds satisfying a $\chi$-type Sobolev inequality, we obtain a lower bound on the volume growth of geodesic balls. We also derive a local logarithmic gradient estimate for positive solutions, assuming that $\mathrm{Ric}_{-}(x)\in L^{\gamma}$ for some $\gamma>\frac{\chi}{\chi-1}$.
Some geometric and topological applications of our main result are also presented in this article (see Theorem 1.10, Theorem 1.11, and Corollory 1.12). In particular, we prove the following. Let $(M,g)$ be a complete noncompact Riemannian manifold of dimension $n\geq 3$ on which the Sobolev inequality (1.8) holds, and assume that $\mathrm{Ric}(x)\geq 0$ outside some geodesic ball $B(o,R_{0})$. Then there exists a positive constant $C(n)$, depending only on $n$, such that if
$\|\mathrm{Ric}_{-}\|_{L^{\frac{n}{2}}}\leq C(n)\,\mathbb{S}_{\frac{n}{n-2}}(M),$
then $(M,g)$ has exactly one end.
2010 Mathematics Subject Classification:
Primary 53C21; Secondary 53C20
1. Introduction
In this paper we are concerned with the following quasi-linear equation
(1.1) $\Delta_{p}v+av^{q}=0$
defined on a complete Riemannian manifold $(M,g)$ which supports a Sobolev inequality, where $p>1$, $a$, $q\in{\mathbb{R}}$ are constants, and the $p$-Laplacian operator is defined as
$\Delta_{p}(v)={\rm div}(|\nabla v|^{p-2}\nabla v).$
In the case $a=1$ and $p=2$, equation (1.1) reduces to the well-known semilinear equation
(1.2) $\Delta v+v^{q}=0,$
commonly referred to as the Lane-Emden equation. This equation arises in various branches of mathematics, such as the prescribed scalar curvature problem (for q=(n+2)/(n2)q=(n+2)/(n-2), cf. e.g. [57, 58]), the scalar field equation (cf. [5]), the stationary solutions to Euler’s equation on $\mathbb{S}^{2}$ (cf. [22, 23]) and has been studied extensively in the last half century (cf. e.g. [1, 7, 19, 21, 26, 29, 33, 34, 35, 36, 46, 53, 50, 71]).
The study on the existence and non-existence of positive solutions to the equation (1.1) and (1.2) is rather subtle. It was proved by Gidas and Spruck in [35] that any nonnegative solutions to (1.2) with 1<q<(n+2)/(n2)1<q<(n+2)/(n-2) on a Riemannian manifold of nonnegative Ricci curvature is zero. Combining the results in [68, 35], we know this result actually holds for $\infty<q<(n+2)/(n-2)$. On the other hand, Ding-Ni proved in [28] that for any $b>0$, there exists a positive solution to (1.2) defined in $\mathbb{R}^{n}$ with $q\geq(n+2)/(n-2)$ such that $\|v\|_{L^{\infty}}=b$. Moreover, Cafarelli-Gidas-Spruck [9] proved that all positive solutions to equation (1.2) in ${\mathbb{R}}^{n}$ with critical power $q=\frac{n+2}{n-2}$ are radial, and can be explicitly written as
(1.3) $u(x)=\left(a+b\lvert x-x_{0}\rvert^{2}\right)^{-\frac{n-2}{2}},\quad\quad n\left(n-2\right)ab=1,$
where $x_{0}\in\mathbb{R}^{n}$, $a>0$ and $b>0$ (cf. [15] for the two dimensional case). We refer to [19, 60] and [18] for recent development of classification results for equation (1.2).
Now, we turn our attention back to the equation (1.1). When $a=0$, this equation becomes the $p$-Laplacian equation
(1.4) $\Delta_{p}v=0,\quad p>1.$
The renowned Cheng-Yau’s logarithm gradient estimate showed that when $p=2$, any solution bounded from above or below to (1.4) is a constant provided the Ricci curvature of the Riemannian manifold is nonnegative (cf. [17]). Kotschwar-Ni [41] proved that any positive $p$-harmonic function on a complete Riemannian manifold with nonnegative sectional curvature is a positive constant. Subsequently, this result was verified to remain correct by Wang and Zhang [65] under the condition of nonnegative Ricci curvature, where the authors applied the Nash-Moser iteration technique and Saloff-Coste’s Sobolev inequality (cf. [56, Theorem 3.1]) to study the gradient estimates of equation (1.4). This gradient estimate was later refined to be sharp by Sung and Wang [62].
When $q\neq p-1$ and $a>0$, the constant $a$ can be absorbed by a dilation transformation, hence equation (1.1) can be reduced to the classical Lane-Emden-Fowler (or Emden-Fowler) equation
(1.5) $\Delta_{p}v+v^{q}=0,$
which appears naturally on fluid mechanics and conformal geometry and has been widely studied in the literature (cf. [5, 6, 7, 38, 51, 52, 59, 61] and the references therein). It was proved by Serrin and Zou in [59] that if $1<p<n$ and $q>0$, then equation (1.5) defined on ${\mathbb{R}}^{n}$ admits no positive solution if and only if
0<q<np/(np)1.0<q<np/(n-p)-1.
By employing the Nash-Moser iteration method, He and the first two named authors of the present paper [38] showed that there is no positive solution of (1.1) defined on a complete Riemannian manifold of nonnegative Ricci curvature with
$a>0\quad\&\quad q<\frac{n+3}{n-1}(p-1)\quad\text{or}\quad a<0\quad\&\quad q>p-1.$
Especially, in the case $p=2$, $a>0$ and $q\in(-\infty,\,\frac{n+2}{n-2})$, Lu [47] established the Cheng-Yau type logarithm gradient estimate for positive solutions to Lane-Emden equation (1.2) on a complete Riemannian manifold with Ricci curvature bounded from below.
Recently, He, Sun and the first named author of this paper [37] showed there is no positive solutions to the subcritical Lane-Emden-Fowler equations(i.e., (1.5) with $\infty<q<\frac{np}{(n-p)_{+}}-1$), over complete Riemannian manifolds with nonnegative Ricci curvature, thereby deriving the optimal Liouville theorems for such equations.
Numerous mathematicians have also studied differential inequalities on a complete manifold $(M,g)$, such as
(1.6) $\Delta_{p}u+u^{q}\leq 0.$
Grigor’yan and Sun [36] and Zhang [71] investigated the uniqueness of a nonnegative solution to this inequality for $p=2$. Notably, they utilized the condition of volume growth instead of relying on nonnegative Ricci curvature. Similar results hold for general $p>1$ (cf. [61]).
Very recently, Ciraolo-Farina-Polvara [18] studied the Liouville theorem for positive solutions to (1.5) defined on a manifold associated with a so-called $\chi$-type Sobolev inequality. In order to introduce their results, we first clarify the definition of this inequality.
Definition 1.1.
Let $\left(M^{n},g\right)$ be an $n$-dimensional Riemannian manifold. We say the $\chi$-type Sobolev inequality holds on $(M,g)$, if $\chi>1$ and there exists a positive constant $\mathbb{S}_{\chi}(M)>0$ such that for any $f\in C^{\infty}_{0}(M,g)$, there holds
(1.7) $\mathbb{S}_{\chi}(M)\left(\int_{M}f^{2\chi}dv\right)^{\frac{1}{\chi}}\leq\int_{M}|\nabla f|^{2}dv.$
It is well known that (1.7) holds for a large class of Cartan–Hadamard manifolds on which Poincaré inequalities are valid; for details, we refer to [39]. We also note that if a $\chi$-type Sobolev inequality holds on $(M,g)$ with $\dim(M)=n\geq 3$, then necessarily $\chi\leq\frac{n}{n-2}$ (see Theorem 2.1). Therefore, throughout the remainder of this article we assume $\chi\in\bigl(1,\frac{n}{n-2}\bigr]$ when $\dim(M)\geq 3$.
Clearly, (1.7) reduces to the classical Sobolev inequality when $\chi=\frac{n}{n-2}$, namely,
(1.8) $\mathbb{S}_{\frac{n}{n-2}}(M)\left(\int_{M}f^{\frac{2n}{n-2}}\,dv\right)^{\frac{n-2}{n}}\leq\int_{M}|\nabla f|^{2}\,dv,$
which holds for a broad class of complete Riemannian manifolds. In recent years, significant effort has been devoted to optimal Sobolev inequalities on Riemannian manifolds; see, for instance, the survey [31] and the references therein.
In the sequel, we will use the notation $\mathrm{Ric}_{-}(\,\cdot\,)$ which is defined as
$\mathrm{Ric}_{-}(x)=\max_{|v|=1,\,v\in T_{x}M}\left\{0,\,-\mathrm{Ric}_{x}(v,v)\right\},\quad\forall x\in M.$
Recently, Ciraolo, Farina, and Polvara (cf. [18]) proved the following results, which may be viewed as a generalization of the classical Gidas–Spruck’s theorem for the semilinear equation $\Delta u+u^{q}=0$ on Riemannian manifolds with nonnegative Ricci curvature to the setting where the Ricci curvature is only bounded in an integral sense.
Theorem 1.2 (Theorem 1.5 in [18]).
Let $(M,g)$ be a complete noncompact Riemannian manifold of dimension $n\geq 3$ on which the $\chi$-type Sobolev inequality holds, and let $u$ be a nonnegative solution to
$\Delta u+u^{q}=0\ \ \ \ \text{in}\ \ M,$
with 1<q<(n+2)/(n2)1<q<(n+2)/(n-2). Assume that
(1.9) $\|\mathrm{Ric}_{-}\|_{L^{\frac{\chi}{\chi-1}}}\leq\frac{1}{48}\left(\frac{n+2}{n-2}-q\right)(q-1)(n-2)\mathbb{S}_{\chi}(M)^{2}.$
Assume also that, for some fixed point $o\in M$, the volume of the geodesic ball $B(o,R)$ satisfies
(1.10) ${\rm{Vol}}(B(o,R))=O\left(R^{2+\frac{8}{q-1}}\right),\ \ \ \text{as}\ \ R\rightarrow\infty.$
Then $u$ is identical to zero.
It was first observed by the first-named author [67, Proposition 2.4], and later independently by Carron (cf. [13]) and Akutagawa (cf. [2]), that if a Riemannian manifold $(M^{n},g)$ has positive Sobolev constant, then the volume of any geodesic ball of radius $R$ satisfies
${\rm{Vol}}(B_{R})\geq CR^{n},$
where the constant $C$ depends only on $n$ and the Sobolev constant.
The first main result of the present article shows that an analogous volume growth estimate also holds for Riemannian manifolds satisfying the $\chi$-type Sobolev inequality. Throughout the paper, we use the abbreviation $B_{r}$ to denote a geodesic ball of radius $r$ centered at an arbitrary point of the manifold. We now state the result precisely as follows.
Theorem 1.3.
Let $(M,g)$ be a complete noncompact Riemannian manifold on which the $\chi$-type Sobolev inequality holds. Then,
$\mathrm{Vol}\left(B_{r}\right)\geq C\left(\chi,\mathbb{S}_{\chi}(M)\right)r^{\frac{2\chi}{\chi-1}}.$
It is worth pointing out that $\frac{2\chi}{\chi-1}=n$ when $\chi=\frac{n}{n-2}$, where $n=\dim(M)$. In this case, we recover the volume estimate obtained in [67]. Moreover, it is easy to see that
$\frac{2\chi}{\chi-1}>2+\frac{8}{q-1}$
whenever $q>4\chi-3$. Consequently, for such $q$ there is a structural incompatibility between the assumptions that a $\chi$-type Sobolev inequality holds and that the volume growth condition (1.10) is satisfied in the theorem of Ciraolo–Farina–Polvara. The main contribution of the present paper is to remove the a priori volume growth assumption from that theorem. We now state our result precisely.
Theorem 1.4.
Let $(M,g)$ be a complete noncompact Riemannian manifold on which the $\chi$-type Sobolev inequality holds. Assume that ${\rm{Vol}}(B(o,R))=O(R^{\beta^{*}})$ for some $\beta^{*}\geq n$ in the case $\chi=\frac{n}{n-2}$, where $B(o,R)\subset M$ is a geodesic ball centered at a fixed point $o\in M$. Then there exists a positive constant $C(n,p,q,\beta^{*})$ depending on $n$, $p$, $q$ and $\beta^{*}$ such that, if
$a=0\quad\text{or}\quad a>0\quad\&\quad q<\frac{n+3}{n-1}(p-1)\quad\text{or}\quad a<0\quad\&\quad q>p-1$
and
$\|\mathrm{Ric}_{-}\|_{L^{\frac{\chi}{\chi-1}}}\leq C(n,p,q,\beta^{*})\mathbb{S}_{\chi}(M),$
then equation (1.1) does not admit any positive solution for $a\neq 0$ and does not admit any non-constant positive solution for $a=0$.
Roughly speaking, we show that if $(M,g)$ satisfies a $\chi$-type Sobolev inequality and has volume growth strictly subexponential, then, provided the $L^{\chi/(\chi-1)}$-norm of $\mathrm{Ric}_{-}$ is bounded by $C^{\prime}\,\mathbb{S}_{\chi}(M)$, where $C^{\prime}$ depends on the growth rate of the volume of geodesic balls, equation (1.1) admits no positive solutions.
Moreover, in the case $p=2$ and $a=1$, we obtain the following conclusion:
Theorem 1.5.
Let $(M,\penalty 10000\ g)$ be a complete noncompact Riemannian manifold on which the $\chi$-type Sobolev inequality holds. Assume that $\mathrm{Vol}\left(B(o,R)\right)=O(R^{\beta^{*}})$ for some $\beta^{*}\geq n$ in the case $\chi=\frac{n}{n-2}$, where $B(o,R)\subset M$ is a geodesic ball centered at a fixed point $o\in M$. Then there exists a positive constant $C(n,q,\beta^{*})$ depending on $n$, $q$ and $\beta^{*}$ such that, if
$q<\frac{n+2}{(n-2)_{+}}$
and
$\|\mathrm{Ric}_{-}\|_{L^{\frac{\chi}{\chi-1}}}\leq C(n,q,\beta^{*})\mathbb{S}_{\chi}(M),$
then Lane-Emden equation (1.2) does not admit any positive solution.
Remark 1.6.
Note that $\frac{\chi}{\chi-1}>\frac{n}{2}$ whenever $\chi\in\bigl(1,\,\frac{n}{n-2}\bigr)$. By the relative volume comparison theorem under integral Ricci curvature bounds due to Petersen and Wei [54], if $\|\mathrm{Ric}_{-}\|_{L^{\frac{\chi}{\chi-1}}}$ is bounded, then
$\mathrm{Vol}\bigl(B(o,R)\bigr)=O\!\left(R^{\frac{2\chi}{\chi-1}}\right)\quad\text{as}\quad R\rightarrow\infty.$
Consequently, the assumption “ $\mathrm{Vol}\bigl(B(o,R)\bigr)=O(R^{\beta^{*}})$ for some $\beta^{*}\geq n$” in Theorem 1.4 and Theorem 1.5 is needed only in the borderline case $\chi=\frac{n}{n-2}$.
As a direct corollary of Theorem 1.4, we obtain the following Liouville type result for $p$ harmonic functions.
Corollary 1.7.
Let $(M,g)$ be a complete noncompact Riemannian manifold on which the Sobolev inequality (1.8) holds true. Assume that $\dim(M)=n\geq 3$ and ${\rm{Vol}}(B(o,R))=O(R^{\beta^{*}})$ for some $\beta^{*}\geq n$ where $B(o,R)$ is a geodesic ball centered at fixed point $o\in M$. Then there exists a positive constant $C(n,p,\beta^{*})$ depending on $n$, $p$ and $\beta^{*}$ such that, if
$\|\mathrm{Ric}_{-}\|_{L^{\frac{n}{2}}}\leq C(n,p,\beta^{*})\mathbb{S}_{\frac{n}{n-2}}(M),$
then there is no nonconstant, positive $p$-harmonic function on $(M,g)$.
Remark 1.8.
Here, we record several remarks on Theorem 1.4 and Theorem 1.5.
  1. (1)
    Our results remove the prior volume growth assumption (1.10).
  2. (2)
    In Theorem 1.2, the restriction $q>1$ is required due to the curvature assumption (1.9). In contrast, our results remain valid for all $q\in(-\infty,1]$.
  3. (3)
    We establish a Liouville theorem for the general $p$-Laplace equation (1.1), where $a$ is an arbitrary constant.
  4. (4)
    In a certain sense, these two theorems may be viewed as effective versions of the classical Liouville results of Gidas–Spruck [35] and Serrin–Zou [59]. To explain this, let $g_{\varepsilon}$ be a family of metrics on $\mathbb{R}^{n}$ such that $g_{\varepsilon}$ agrees with the Euclidean metric outside a compact set and $g_{\varepsilon}\to g_{\mathrm{Euc}}$ in $C^{2}$ as $\varepsilon\to 0$. The Positive Mass Theorem implies that $g_{\varepsilon}$ cannot have nonnegative Ricci curvature. Since the Sobolev constant depends only on the $C^{0}$-geometry of the metric, our results show that the classical Liouville property still holds for such a deformation provided that
    $\bigl\|\mathrm{Ric}^{g_{\varepsilon}}_{-}\bigr\|_{L^{\frac{n}{2}}}\leq C(n,p,q)\,\mathbb{S}_{\frac{n}{n-2}}(\mathbb{R}^{n},g_{\varepsilon}).$
Assume that $\mathrm{Ric}_{-}(x)\in L^{\gamma}(B_{1})$ for some $\gamma>\frac{\chi}{\chi-1}$. By the relative volume comparison theorem under integral Ricci curvature bounds due to Petersen and Wei (cf. [54]), we obtain the following local gradient estimate for solutions of (1.1).
Theorem 1.9.
Let $(M,g)$ be a complete noncompact Riemannian manifold with $\dim(M)\geq 3$ on which the $\chi$-type Sobolev inequality holds. Assume $v$ is a positive solution to (1.1) on $B_{1}\subset M$. Then, when
$a=0\quad\text{or}\quad a>0\quad\&\quad q<\frac{n+3}{n-1}(p-1)\quad\text{or}\quad a<0\quad\&\quad q>p-1,$
the following gradient estimate holds true
(1.11) $\sup_{B_{1/2}}\frac{|\nabla v|^{2}}{v^{2}}\leq C\left(p,q,\mathbb{S}_{\chi}(M),\gamma,\|\mathrm{Ric}_{-}\|_{L^{\gamma}(B_{1})}\right),$
where $\gamma$ is any number greater than $\chi/(\chi-1)$. In particular, if $\|\mathrm{Ric}_{-}\|_{L^{\gamma}(M)}\leq\Lambda$, then the following global gradient estimate holds
$\frac{|\nabla v|^{2}}{v^{2}}\leq C\left(p,q,\mathbb{S}_{\chi}(M),\gamma,\Lambda\right).$
In [55, Theorem 1.2 ], Petersen and Wei proved that if a manifold has positive Sobolev constant and $\mathrm{Ric}_{-}\in L^{p}$ for some $p>\frac{n}{2}$, then any positive harmonic function $v$ on $B_{1}$ satisfies the local gradient estimate
$\sup_{B_{1/2}}|\nabla v|\leq C\sup_{B_{1}}v.$
Therefore, Theorem 1.9 both improves and extends the result of Petersen and Wei.
Now, we turn to recalling some topological features of noncompact complete manifolds, with a focus on estimates for the number of ends. Notice that “an end $E$ of a manifold $M$ is an unbounded component of the complement of some compact subset $D$ of $M$. In this case, we say that $E$ is an end corresponding to DD”.
For a noncompact complete manifold whose Ricci curvature is nonnegative outside a compact set, Cai [10] proved that such a manifold has only finitely many ends. Later, in 1995, Cai, Colding, and Yang [11] studied a gap phenomenon for the ends of this class of manifolds. More precisely, they proved the following theorem:
Given $n>0$, there exists an $\epsilon(n)>0$ such that for all pointed open complete manifolds $(M^{n},o)$ with Ricci curvature bounded from below by $(n-1)\Lambda^{2}$ and nonnegative outside the ball $B(o,a)$, if $\Lambda a<\epsilon(n)$, then $M^{n}$ has at most two ends.
For more results related to this topic, we refer to [20, 69] and references therein.
On the other hand, using harmonic function theory to study the number of ends has a long history (see for example, [12, 30, 42]). Especially, if the Sobolev constant of $(M,g)$ is positive and $\dim(M)\geq 3$, the first named author has ever used harmonic function theory to prove that such a manifold is of finitely many ends and the dimension of linear space spanned by bounded harmonic functions on $(M,g)$ equals the number of ends (cf. [67]).
By the volume estimate established in Theorem 1.3 we can employ the method adopted in [30] and [67] to show the following theorem.
Theorem 1.10.
Let $(M,g)$ be a complete noncompact Riemannian manifold with Sobolev constant $\mathbb{S}_{\chi}(M)>0$. If $(M,g)$ has at least $k$ ends, then $\dim(H^{\infty}(M))\geq k$.
As a geometric application of Corollory 1.7 and Theorem 1.10, we obtain the following gap theorem for ends.
Theorem 1.11.
Let $(M,g)$ be a complete noncompact Riemannian manifold on which the Sobolev inequality (1.8) holds. Assume that $\dim(M)=n\geq 3$ and ${\rm{Vol}}(B(o,R))=O(R^{\beta^{*}})$ for some $\beta^{*}\geq n$ where $B(o,R)$ is a geodesic ball centered at fixed point $o\in M$.Then there exists a positive constant $C(n,\beta^{*})$ depending only on $n$ and $\beta^{*}$ such that, if
$\|\mathrm{Ric}_{-}\|_{L^{\frac{n}{2}}}\leq C(n,\beta^{*})\mathbb{S}_{\frac{n}{n-2}}(M),$
then $(M,g)$ is of a unique end.
As a direct corollary of the above theorem, we have the following:
Corollary 1.12.
Let $(M,g)$ be a complete noncompact Riemannian manifold with $\dim(M)=n\geq 3$ on which the Sobolev inequality (1.8) holds true. Assume Ricci curvature is nonnegative outside some compact set, or more generally, $\mathrm{Vol}\left(B_{r}\right)=O(r^{n})$ for any $r\to\infty$. Then there exists a positive constant $C(n)$ depending only on $n$ such that, if
$\|\mathrm{Ric}_{-}\|_{L^{\frac{n}{2}}}\leq C(n)\mathbb{S}_{\frac{n}{n-2}}(M),$
then $(M,g)$ is of a unique end. In particular, if $(M,g)$ is a complete Riemannian manifold with $\dim(M)=n\geq 3$ and nonnegative Ricci curvature on which the Sobolev inequality (1.8) holds true, then $(M,g)$ has only an end.
Combining Theorem 1.11 and Cai-Colding-Yang’s result, we would like to ask the following problem: whether or not there exists a positive constant $\epsilon(n)$ depending on $n$ such that any noncompact complete Riemannian manifold with nonnegative Ricci curvature outside a compact set is of at most two ends if
$\|\mathrm{Ric}_{-}\|_{L^{\frac{n}{2}}}\leq\epsilon(n);$
or more generally, does there exist $\epsilon(n)>0$ such that any noncompact complete Riemannian manifold has at most two ends whenever
$\|\mathrm{Ric}_{-}\|_{L^{\frac{n}{2}}}\leq\epsilon(n)?$
The rest of the paper is organized as follows. In Section 2, we derive a volume estimate for geodesic balls $B_{R}$ in $(M,g)$ under the assumption that $(M,g)$ satisfies a $\chi$-type Sobolev inequality. Section 3 is devoted to a careful estimate of ${\mathscr{L}}\!\left(|\nabla\log v|^{2\alpha}\right)$ (the operator $\mathscr{L}$} is defined explicitly in (3.3)). Our proofs show that, by choosing a suitable parameter $\alpha$, one can obtain effective integral estimates for the gradient of positive solutions to (1.1). In particular, under the assumptions of Theorem 1.4 or Theorem 1.9, we obtain an $L^{\theta\chi}$ bound for $|\nabla\log v|$. Crucially, this bound depends only on the radius and the volume of the geodesic ball. With this integral bound in hand, we then apply the Nash–Moser iteration scheme to complete the proof of Theorem 1.9. In Section 4, we proceed to prove Theorem 1.5 by choosing appropriate auxiliary functions. Finally, in Section 5 we prove Theorem 1.10 and Theorem 1.11.
2. Volume estimate
In this section, we shall provide two different proofs of Theorem 1.3. One of the two proofs is to make use of the Nash-Moser iteration initially developed by Wang in [67], the other is a direct iteration of volume of the geodesic ball as in [39].
First, we will show the following result.
Theorem 2.1.
Let $(M^{n},g)$ be a complete noncompact Riemannian manifold on which the $\chi$-type Sobolev inequality holds with $n\geq 3$, then $\chi\leq n/(n-2)$.
Proof.
Fix a point $o\in M$ and fix some $r>0$. Define
$u(x)=\left\{\begin{aligned} r-d_{g}(x,0)&,\ \ \ \ \text{if}\ \ d_{g}(x,o)\leq r,\\ 0&,\ \ \ \ \text{if}\ \ d_{g}(x,0)\geq r,\end{aligned}\right$.
where $d_{g}(\cdot,\cdot)$ is the distance function on $(M,g)$. Obviously, $u\in W_{0}^{1,2}(M,g)$. Notice that
$\int_{M}|\nabla u|^{2}=\mathrm{Vol}\left(B_{r}(o)\right).$
On the other hand, we know that
$\mathrm{Vol}\left(B_{r}(o)\right)=\omega_{n}r^{n}(1+o(1)),\quad\mathrm{and}\quad\mathrm{Area}\left(\partial B_{r}(o)\right)=n\omega_{n}r^{n-1}(1+o(1))\ \ \ \text{as}\ r\rightarrow 0.$
Direct calculation shows that
$\left(\int_{M}u^{2\chi}\right)^{\frac{1}{\chi}}=\left(n\omega_{n}\frac{\Gamma(n)\Gamma(2\chi+1)}{\Gamma(n+2\chi+1)}\right)^{\frac{1}{\chi}}r^{2+\frac{n}{\chi}}(1+o(1))\ \ \text{as}\ r\rightarrow 0.$
From the definition of the $\chi$-type Sobolev inequality, there must hold
$n\leq 2+\frac{n}{\chi}.$
Hence, $\chi\leq n/(n-2)$. ∎
Now, we will establish a local maximum principle for subharmonic functions on Riemannian manifolds on which the $\chi$-type Sobolev inequality holds via the classical Nash-Moser iteration. Then, we will see later that Theorem 1.3 is a direct corollary of this local maximum principle.
Lemma 2.2.
Let $(M,g)$ be a complete noncompact Riemannian manifold on which the $\chi$-type Sobolev inequality holds. Assume $u\in W^{1,2}\left(B_{r}\right)$ satisfying
$\Delta u\geq 0,$
in the weak sense, i.e.,
$\int_{B_{r}}\langle\nabla u,\nabla\varphi\rangle\leq 0,\ \ \text{for any}\ \ 0\leq\varphi\in C^{\infty}_{0}\left(B_{r}\right).$
Then, for any $s>0$ and $0<\theta<1$, there holds
(2.1) $\sup_{B_{\theta r}}u\leq C\left(\chi,s,\theta,\mathbb{S}_{\chi}(M)^{-1}\right)r^{-\frac{2\chi}{s(\chi-1)}}\left(\int_{B_{r}}(u^{+})^{s}\right)^{1/s}.$
Proof.
Since $u^{+}$ is also the subsolution, without loss of generality we assume $u\geq 0$. We first prove this lemma in the case $s\geq 2$. By integration by part, we know
$\int_{B_{r}}\langle\nabla u,\nabla\varphi\rangle\leq 0,\ \ \text{for any}\ \ 0\leq\varphi\in W^{1,2}_{0}\left(B_{r}\right).$
For any $\eta(x)\in C_{0}^{\infty}\left(B_{r}\right)$, substituting $\varphi=\eta^{2}u^{s-1}$ into the above inequality yields
(2.2) $(s-1)\int_{B_{r}}|\nabla u|^{2}u^{s-2}\eta^{2}\leq-2\int_{B_{r}}\eta u^{s-1}\langle\nabla u,\nabla\eta\rangle.$
By Young’s inequality, we deduce that
(2.3) $\int_{B_{r}}|\nabla u|^{2}u^{s-2}\eta^{2}\leq\frac{4}{(s-1)^{2}}\int_{B_{r}}u^{s}|\nabla\eta|^{2}.$
Note that
$u^{s-2}|\nabla u|^{2}=\frac{4}{s^{2}}|\nabla u^{\frac{s}{2}}|^{2}.$
We know
$|\nabla(\eta u^{s/2})|^{2}=\eta^{2}|\nabla u^{\frac{s}{2}}|^{2}+u^{s}|\nabla\eta|^{2}+s\eta u^{s-1}\langle\nabla u,\nabla\eta\rangle.$
This implies
(2.4) $\int_{B_{r}}|\nabla(\eta u^{s/2})|^{2}$ $=$ $\int\eta^{2}|\nabla u^{\frac{s}{2}}|^{2}+\int u^{s}|\nabla\eta|^{2}+s\int\eta u^{s-1}\langle\nabla u,\nabla\eta\rangle$
$\leq$ $\frac{s^{2}}{(s-1)^{2}}\int u^{s}|\nabla\eta|^{2}+\int u^{s}|\nabla\eta|^{2}+\int u^{s}|\nabla\eta|^{2}+\frac{s^{2}}{4}\int u^{s-2}|\nabla u|^{2}\eta^{2}$
$\leq$ $2\left(1+\frac{s^{2}}{(s-1)^{2}}\right)\int u^{s}|\nabla\eta|^{2}$
$\leq$ $10\int u^{s}|\nabla\eta|^{2}.$
By Sobolev inequality, there holds
$\mathbb{S}_{\chi}(M)\left(\int_{B_{r}}(\eta u^{s/2})^{2\chi}\right)^{\frac{1}{\chi}}\leq\int_{B_{r}}|\nabla(\eta u^{s/2})|^{2}.$
Combining the above inequality with (2.4), we arrive at
(2.5) $\left(\int_{B_{r}}(\eta u^{s/2})^{2\chi}\right)^{\frac{1}{\chi}}\leq 10\mathbb{S}_{\chi}(M)^{-1}\int u^{s}|\nabla\eta|^{2}.$
For some positive number $\theta\in(0,1)$, let
$r_{k}=r\left(\theta+\frac{1-\theta}{2^{k}}\right),\ \ \ k=0,1,2,\cdots$
and choose $\eta_{k}\in C_{0}^{\infty}\left(B_{r_{k}}\right)$ such that $\eta_{k}\equiv 1$ on $B_{r_{k+1}}$ and
(2.6) $|\nabla\eta_{k}|\leq\frac{2}{r_{k}-r_{k+1}}=\frac{2^{k+2}}{(1-\theta)r}.$
Let
$s_{k}=s\chi^{k}.$
Substituting $\eta\triangleq\eta_{k}$, $s\triangleq s_{k}$ and (2.6) into (2.5), we obtain
(2.7) $\|u\|_{L^{s_{k+1}}\left(B_{r_{k+1}}\right)}\leq\left\{\frac{160\times 4^{k}}{\mathbb{S}_{\chi}(M)(1-\theta)^{2}r^{2}}\right\}^{\frac{1}{s_{k}}}\|u\|_{L^{s_{k}}\left(B_{r_{k}}\right)}.$
By iteration, we derive
(2.8) $\|u\|_{L^{s_{k+1}}\left(B_{r_{k+1}}\right)}\leq 4^{\sum_{i=0}^{k}\frac{i}{s_{i}}}\left\{\frac{160}{\mathbb{S}_{\chi}(M)(1-\theta)^{2}r^{2}}\right\}^{\sum_{i=0}^{k}\frac{1}{s_{i}}}\|u\|_{L^{s}\left(B_{r}\right)}.$
Since
$\sum_{i=0}^{\infty}\frac{1}{s_{i}}=\frac{\chi}{s(\chi-1)}\quad\mbox{and}\quad\sum_{i=0}^{\infty}\frac{i}{s_{i}}=\frac{\chi}{s(\chi-1)^{2}}.$
Letting $k\rightarrow\infty$ in (2.8) yields
(2.9) $\|u\|_{L^{\infty}\left(B_{\theta r}\right)}\leq C\left(\chi,s,\mathbb{S}_{\chi}(M)^{-1}\right)[(1-\theta)r]^{-\frac{2\chi}{s(\chi-1)}}\|u\|_{L^{s}\left(B_{r}\right)}.$
Next, we prove Lemma 2.2 when $s\in(0,2)$. By letting $s=2$ in (2.9), we obtain
$\sup_{B_{\theta r}}u\leq C\left(\chi,\mathbb{S}_{\chi}(M)^{-1}\right)[(1-\theta)r]^{-\frac{\chi}{\chi-1}}\|u\|_{L^{2}\left(B_{r}\right)}.$
Hence, for $0<s<2$, there holds true
$\sup_{B_{\theta r}}u\leq C\left(\chi,\mathbb{S}_{\chi}(M)^{-1}\right)[(1-\theta)r]^{-\frac{\chi}{\chi-1}}\left(\sup_{B_{r}}u\right)^{1-\frac{s}{2}}\left(\int_{B_{r}}u^{s}\right)^{\frac{1}{2}}.$
By Young’s inequality, we deduce that
(2.10) $\sup_{B_{\theta r}}u\leq\frac{2-s}{2}\sup_{B_{r}}u+\frac{s}{2}C\left(\chi,\mathbb{S}_{\chi}(M)^{-1}\right)^{\frac{2}{s}}[(1-\theta)r]^{-\frac{2\chi}{s(\chi-1)}}\|u\|_{L^{s}\left(B_{r}\right)}.$
Let $\widetilde{s}=\theta r$, $t=r$, $\psi(s)=\sup_{B_{\theta r}}u$ and $\psi(t)=\sup_{B_{r}}u$. Then (2.10) can be rewritten as
$\psi(\widetilde{s})\leq\frac{2-s}{2}\psi(t)+C\left(\chi,\mathbb{S}_{\chi}(M)^{-1}\right)(t-\widetilde{s})^{-\frac{2\chi}{s(\chi-1)}}\|u\|_{L^{s}\left(B_{r}\right)}.$
By Lemma 2.3 below, we conclude that for $s\in(0,\,2)$, inequality (2.9) also holds. Thus, we complete the proof. ∎
Lemma 2.3 (cf. [16]).
Let $f(t)\geq 0$, $t\in[\tau_{0},\tau_{1}]$ with $\tau_{0}\geq 0$. Suppose for $\tau_{0}\leq t<\widetilde{s}\leq\tau_{1}$,
$f(t)\leq\theta f(\widetilde{s})+\frac{A}{(\widetilde{s}-t)^{\alpha}}+B$
for some $\theta\in[0,1)$. Then for any $\tau_{0}\leq t<\widetilde{s}\leq\tau_{1}$, there holds
$f(t)\leq c(\alpha,\theta)\left(\frac{A}{(\widetilde{s}-t)^{\alpha}}+B\right).$
Proof of Theorem 1.3(Method 1):.
It is easy to see that Theorem 1.3 can be directly deduced from Lemma 2.2 by letting $u=1$, $s=1$ and $\theta=\frac{1}{2}$ in (2.1). ∎
Next, we shall prove Theorem 1.3 via a direct iteration of the volume of the geodesic balls as in [39].
Proof of Theorem 1.3:.
Recall that the $\chi$-type Sobolev inequality tells us that for any $u\in W^{1,2}_{0}(M)$, there holds
$\mathbb{S}_{\chi}(M)\left(\int_{M}u^{2\chi}dv\right)^{\frac{1}{\chi}}\leq\int_{M}|\nabla u|^{2}.$
Now, let $r>0$ and $x$ be some point of $M$, and let $u\in W^{1,2}_{0}(M)$ be such that $u=0$ on $M\backslash B_{x}(r)$. By Hölder’s inequality, we have
$\left(\int_{M}u^{2}dv\right)^{\frac{1}{2}}\leq\left(\int_{M}u^{2\chi}dv\right)^{\frac{1}{2\chi}}{\rm{Vol}}\left(B_{x}(r)\right)^{\frac{\chi-1}{2\chi}}.$
Hence,
(2.11) $\frac{\left(\int_{M}|\nabla u|^{2}dv\right)^{\frac{1}{2}}}{\left(\int_{M}u^{2}dv\right)^{\frac{1}{2}}}$ $\geq$ $\sqrt{\mathbb{S}_{\chi}(M)}\frac{\left(\int_{M}u^{2\chi}dv\right)^{\frac{1}{2\chi}}}{\left(\int_{M}u^{2}dv\right)^{\frac{1}{2}}}$
$\geq$ $\frac{\sqrt{\mathbb{S}_{\chi}(M)}}{{\rm{Vol}}\left(B_{x}(r)\right)^{\frac{\chi-1}{2\chi}}}.$
From now on, let
$u(y)=\left\{\begin{aligned} r-d_{g}(x,y)&,\ \ \ \ \text{if}\ \ d_{g}(x,y)\leq r,\\ 0&,\ \ \ \ \text{if}\ \ d_{g}(x,y)\geq r,\end{aligned}\right$.
where $d_{g}(\cdot,\cdot)$ is the distance function on $(M,g)$. Obviously, $u$ is Lipschitz and $u=0$ on $M\backslash B_{x}(r)$. Substituting $u$ into (2.11) yields
$\frac{\int_{M}|\nabla u|^{2}dv}{\int_{M}u^{2}dv}$ $=$ $\frac{{\rm{Vol}}\left(B_{x}(r)\right)}{\int_{B_{x}(r)}u^{2}dv}$
$\geq$ $\frac{\mathbb{S}_{\chi}(M)}{{\rm{Vol}}\left(B_{x}(r)\right)^{\frac{\chi-1}{\chi}}}.$
Note that
$\int_{B_{x}(r)}u^{2}dv\geq\int_{B_{x}\left(r/2\right)}u^{2}dv,$
and
$\int_{B_{x}\left(r/2\right)}u^{2}dv\geq\frac{r^{2}}{2^{2}}{\rm{Vol}}\left(B_{x}\left(\frac{r}{2}\right)\right).$
It is easy to see that
$\frac{\mathbb{S}_{\chi}(M)}{{\rm{Vol}}\left(B_{x}(r)\right)^{\frac{\chi-1}{\chi}}}\leq\frac{{\rm{Vol}}\left(B_{x}(r)\right)}{\int_{B_{x}\left(r/2\right)}u^{2}dv}\leq\frac{2^{2}{\rm{Vol}}\left(B_{x}(r)\right)}{r^{2}{\rm{Vol}}\left(B_{x}\left(\frac{r}{2}\right)\right)}.$
Hence, we conclude from the above inequality that
${\rm{Vol}}\left(B_{x}(r)\right)\geq\left(\frac{r\sqrt{\mathbb{S}_{\chi}(M)}}{2}\right)^{\frac{2\chi}{2\chi-1}}{\rm{Vol}}\left(B_{x}\left(\frac{r}{2}\right)\right)^{\frac{\chi}{2\chi-1}}.$
Hence, for any $m\in\mathbb{N}$, there holds
${\rm{Vol}}\left(B_{x}\left(\frac{r}{2^{m}}\right)\right)\geq\left(r\sqrt{\mathbb{S}_{\chi}(M)}\right)^{\frac{2\chi}{2\chi-1}}{2}^{-\frac{2(m+1)\chi}{2\chi-1}}{\rm{Vol}}\left(B_{x}\left(\frac{r}{2^{m+1}}\right)\right)^{\frac{\chi}{2\chi-1}}.$
By induction, we then arrive at
(2.12) ${\rm{Vol}}\left(B_{x}(r)\right)\geq\left(r\sqrt{\mathbb{S}_{\chi}(M)}\right)^{2\alpha(m)}{2}^{-2\beta(m)}{\rm{Vol}}\left(B_{x}\left(\frac{r}{2^{m}}\right)\right)^{\gamma(m)},$
where
$\alpha(m)=\sum^{m}_{i=1}\left(\frac{\chi}{2\chi-1}\right)^{i},\ \ \beta(m)=\sum^{m}_{i=1}i\left(\frac{\chi}{2\chi-1}\right)^{i},$
and
$\gamma(m)=\left(\frac{\chi}{2\chi-1}\right)^{m}.$
Direct computation shows that
$\lim_{m\rightarrow\infty}\alpha(m)=\frac{\chi}{\chi-1}\quad\quad\text{and}\quad\quad\lim_{m\rightarrow\infty}\beta(m)=\frac{\chi(2\chi-1)}{(\chi-1)^{2}}.$
On the other hand, it’s well known that the volume of geodesic ball with radius $r$ has the following expansion (cf. [32])
${\rm{Vol}}\left(B_{x}(r)\right)=b_{n}r^{n}\left(1-\frac{\mathrm{R}_{g}(x)}{6(n+2)}r^{2}+o\left(r^{2}\right)\right),$
where $\mathrm{R}_{g}(x)$ denotes the scalar curvature of $(M,g)$ at $x$ and $b_{n}$ is the volume of the Euclidean ball of radius one. Hence,
$\lim_{m\rightarrow\infty}{\rm{Vol}}\left(B_{x}\left(\frac{r}{2^{m}}\right)\right)^{\gamma(m)}=1.$
By letting $m\rightarrow\infty$, we obtain
${\rm{Vol}}\left(B_{x}(r)\right)\geq C\left(\chi,\mathbb{S}_{\chi}(M)\right)r^{\frac{2\chi}{\chi-1}}.$
Thus we complete the proof of Theorem 1.3. ∎
3. $p$-Laplace case: Proof of Theorem 1.4 and Theorem 1.9
3.A. Linearization operator $\mathscr{L}$} of $p$-Laplacian.
Recall that the $p$-Laplace operator is defined as
(3.1) $\Delta_{p}u={\rm div}\left(|\nabla u|^{p-2}\nabla u\right).$
The solution of $p$-Laplace equation $\Delta_{p}u=0$, usually called $p$-harmonic function, is the critical point of the energy functional
$E(u)=\int_{M}|\nabla u|^{p}.$
From the definition, we see that a $2$-harmonic function is just a usual harmonic function.
Definition 3.1.
$v$ is said to be a (weak) solution of equation (1.1) on a region $\Omega\subset M$, if $v\in L^{\infty}_{loc}(\Omega)\cap W^{1,p}_{loc}(\Omega)$ and for all $\psi\in W^{1,p}_{0}(\Omega)$, we have
$-\int_{\Omega}|\nabla v|^{p-2}\langle\nabla v,\nabla\psi\rangle+\int_{\Omega}av^{q}\psi=0.$
From now on, we always assume that $v\in W_{loc}^{1,p}\left(\Omega\right)\cap L_{loc}^{\infty}\left(\Omega\right)$ is a weak and positive solution of the equation (1.1). We denote
$\Omega_{cr}=\{x\in\Omega:\nabla v(x)=0\}.$
According to Theorem 1.4 in [3] and the classical regularity theory (for example, see [27, 59, 63, 64]), we know that
$v\in C_{loc}^{1,\beta}(\Omega)\cap W^{2,2}_{loc}(\Omega\setminus\Omega_{cr})\quad\mbox{and}\quad v\in C_{loc}^{\infty}(\Omega^{c}_{cr}).$
On the other hand, it is easy to see from (3.1) that the linearization operator $\mathscr{L}$} of the $p$-Laplace operator is
${\mathscr{L}}(\psi)={\rm div}\left(|\nabla u|^{p-2}A(\nabla\psi)\right),$
where
$A(\nabla\psi)=\nabla\psi+(p-2)|\nabla u|^{-2}\langle\nabla\psi,\nabla u\rangle\nabla u.$
Now, let $v$ be a positive solution to equation (1.1). By a logarithmic transformation
$u=-(p-1)\log v,$
equation (1.1) becomes
(3.2) $\Delta_{p}u-|\nabla u|^{p}-be^{cu}=0,$
where
$b=a(p-1)^{p-1},\quad c=\frac{p-q-1}{p-1}.$
Denote $f=|\nabla u|^{2}$. Then the linearization operator $\mathscr{L}$} of the $p$-Laplace operator can be rewritten as
(3.3) ${\mathscr{L}}(\psi)={\rm div}\left(f^{p/2-1}A(\nabla\psi)\right),$
with
(3.4) $A(\nabla\psi)=\nabla\psi+(p-2)f^{-1}\langle\nabla\psi,\nabla u\rangle\nabla u.$
Next, we calculate the explicit expression ${\mathscr{L}}(f^{\alpha})$ for any $\alpha>0$ that will play a key role in our proof.
Lemma 3.2.
For any $\alpha>0$, the equality
(3.5) $\begin{split}{\mathscr{L}}(f^{\alpha})=&\alpha\left(\alpha+\frac{p}{2}-2\right)f^{\alpha+\frac{p}{2}-3}|\nabla f|^{2}+2\alpha f^{\alpha+\frac{p}{2}-2}\left(|\nabla\nabla u|^{2}+\mathrm{Ric}(\nabla u,\nabla u)\right)\\ &+\alpha(p-2)(\alpha-1)f^{\alpha+\frac{p}{2}-4}\langle\nabla f,\nabla u\rangle^{2}+2\alpha f^{\alpha-1}\langle\nabla\Delta_{p}u,\nabla u\rangle\end{split}$
holds point-wisely in $\{x:f(x)>0\}$.
Proof.
By the definition of AA in (3.4), we have
$A\Big(\nabla(f^{\alpha})\Big)=\alpha f^{\alpha-1}\nabla f+\alpha(p-2)f^{\alpha-2}\langle\nabla f,\nabla u\rangle\nabla u=\alpha f^{\alpha-1}A(\nabla f).$
Hence
${\mathscr{L}}(f^{\alpha})$ $=\alpha\text{div}\Big(f^{\alpha-1}f^{\frac{p}{2}-1}A(\nabla f)\Big)=\alpha\Big\langle\nabla(f^{\alpha-1}),f^{\frac{p}{2}-1}A(\nabla f)\Big\rangle+\alpha f^{\alpha-1}\mathcal{L}(f).$
A straightforward computation shows that
(3.6) $\alpha\Big\langle\nabla(f^{\alpha-1}),f^{\frac{p}{2}-1}A(\nabla f)\Big\rangle=$ $\Big\langle\alpha(\alpha-1)f^{\alpha-2}\nabla f,f^{\frac{p}{2}-1}\nabla f+(p-2)f^{\frac{p}{2}-2}\langle\nabla f,\nabla u\rangle\nabla u\Big\rangle,$
and
(3.7) $\begin{split}\alpha f^{\alpha-1}{\mathscr{L}}(f)=&\alpha f^{\alpha-1}\Big(\left(\frac{p}{2}-1\right)f^{\frac{p}{2}-2}|\nabla f|^{2}+f^{\frac{p}{2}-1}\Delta f+(p-2)\left(\frac{p}{2}-2\right)f^{\frac{p}{2}-3}\langle\nabla f,\nabla u\rangle^{2}\\ &+(p-2)f^{\frac{p}{2}-2}\langle\nabla\langle\nabla f,\nabla u\rangle,\nabla u\rangle+(p-2)f^{\frac{p}{2}-2}\langle\nabla f,\nabla u\rangle\Delta u\Big).\end{split}$
Combining (3.6) and (3.7) yields
(3.8) $\begin{split}{\mathscr{L}}(f^{\alpha})=&\alpha\left(\alpha+\frac{p}{2}-2\right)f^{\alpha+\frac{p}{2}-3}|\nabla f|^{2}+\alpha f^{\alpha+\frac{p}{2}-2}\Delta f\\ &+\alpha(p-2)\left(\alpha+\frac{p}{2}-3\right)f^{\alpha+\frac{p}{2}-4}\langle\nabla f,\nabla u\rangle^{2}\\ &+\alpha(p-2)f^{\alpha+\frac{p}{2}-3}\langle\nabla\langle\nabla f,\nabla u\rangle,\nabla u\rangle+\alpha(p-2)f^{\alpha+\frac{p}{2}-3}\langle\nabla f,\nabla u\rangle\Delta u.\end{split}$
Notice that, by the definition of the $p$-Laplacian we have
$\langle\nabla\Delta_{p}u,\nabla u\rangle$ $=\left(\frac{p}{2}-1\right)\left(\frac{p}{2}-2\right)f^{\frac{p}{2}-3}\langle\nabla f,\nabla u\rangle^{2}+\left(\frac{p}{2}-1\right)f^{\frac{p}{2}-2}\langle\nabla\langle\nabla f,\nabla u\rangle,\nabla u\rangle$
$\quad+\left(\frac{p}{2}-1\right)f^{\frac{p}{2}-2}\langle\nabla f,\nabla u\rangle\Delta u+f^{\frac{p}{2}-1}\langle\nabla\Delta u,\nabla u\rangle.$
Hence, the last term of the right hand side of (3.8) can be rewritten as
(3.9) $\begin{split}\alpha(p-2)f^{\alpha+\frac{p}{2}-3}\langle\nabla f,\nabla u\rangle\Delta u&=2\alpha f^{\alpha-1}\langle\nabla\Delta_{p}u,\nabla u\rangle-2\alpha f^{\alpha+\frac{p}{2}-2}\langle\nabla\Delta u,\nabla u\rangle\\ &\quad-\alpha(p-2)\left(\frac{p}{2}-2\right)f^{\alpha+\frac{p}{2}-4}\langle\nabla f,\nabla u\rangle^{2}\\ &\quad-\alpha(p-2)f^{\alpha+\frac{p}{2}-3}\langle\nabla\langle\nabla f,\nabla u\rangle,\nabla u\rangle.\end{split}$
Moreover, the Bochner formula tells us that
$\langle\nabla\Delta u,\nabla u\rangle=\frac{1}{2}\Delta f-|\nabla\nabla u|^{2}-\mathrm{Ric}(\nabla u,\nabla u).$
By substituting the above and (3.9) into (3.8), we finally arrive at
${\mathscr{L}}(f^{\alpha})=$ $\alpha\left(\alpha+\frac{p}{2}-2\right)f^{\alpha+\frac{p}{2}-3}|\nabla f|^{2}+2\alpha f^{\alpha+\frac{p}{2}-2}\left(|\nabla\nabla u|^{2}+\mathrm{Ric}(\nabla u,\nabla u)\right)$
$+\alpha(p-2)(\alpha-1)f^{\alpha+\frac{p}{2}-4}\langle\nabla f,\nabla u\rangle^{2}+2\alpha f^{\alpha-1}\langle\nabla\Delta_{p}u,\nabla u\rangle.$
We finish the proof of the lemma. ∎
3.B. Precise estimate of $\mathscr{L}$}.
In this subsection, we shall prove a precise estimate for ${\mathscr{L}}(f^{\alpha})$ when $v$ is a positive solution to equation (1.1).
Lemma 3.3.
Let $u$ be a solution of equation (3.2) on $(M,g)$. Denote
$f=|\nabla u|^{2}\ \ \ \text{and}\ \ \ a_{1}=\left|p-\frac{2(p-1)}{n-1}\right|.$
Then the following holds point-wisely in $\{x\in M:f(x)>0\}$:
(3.10) ${\mathscr{L}}(f^{\alpha})\geq 2\alpha f^{\alpha+\frac{p}{2}-2}\left(\beta_{n,p,q,\alpha}f^{2}-\mathrm{Ric}_{-}f-\frac{a_{1}}{2}f^{\frac{1}{2}}|\nabla f|\right),$
provided
  1. (1)
    $\alpha\in[1,\infty)$ and $\beta_{n,p,q,\alpha}=1/(n-1)$ when
    $a\left(\frac{n+1}{n-1}-\frac{q}{p-1}\right)\geq 0,$
  2. (2)
    $\alpha>\alpha_{0}$, where
    $\alpha_{0}(n,p,q)=\frac{\frac{4}{n-1}+(p-n)\left(\frac{n+1}{n-1}-\frac{q}{p-1}\right)^{2}}{2\left(\frac{4}{n-1}-(n-1)(\frac{n+1}{n-1}-\frac{q}{p-1})^{2}\right)},$
    and
    $\beta_{n,p,q,\alpha}=\frac{1}{n-1}-\left(\frac{n+1}{n-1}-\frac{q}{p-1}\right)^{2}\frac{(2\alpha-1)(n-1)+p-1}{4(2\alpha-1)}>0$
    when
    $p-1<q<\frac{n+3}{n-1}(p-1).$
Proof.
Let $\{e_{1},e_{2},\ldots,e_{n}\}$ be an orthonormal frame of $TM$ on a domain of $\{x\in M:f(x)>0\}$ such that $\nabla u=|\nabla u|e_{1}$. Under this frame, there holds
$u_{1}=|\nabla u|=f^{1/2},\quad\text{and}\quad u_{i}=0\quad\text{for}\ \ 2\leq i\leq n.$
Moreover, notice that
$2fu_{11}$ $=$ $2|\nabla u|^{2}\nabla^{2}u(e_{1},e_{1})$
$=$ $\langle\nabla u,\nabla f\rangle,$
and
$|\nabla f|^{2}$ $=$ $\sum_{i=1}^{n}|2u_{1}u_{1i}|^{2}=4f\sum_{i=1}^{n}u_{1i}^{2}.$
Hence,
(3.11) $u_{11}=\frac{1}{2}f^{-1}\langle\nabla u,\nabla f\rangle\ \ \ \text{and}\ \ \ \frac{|\nabla f|^{2}}{f}=4\sum_{i=1}^{n}u_{1i}^{2}.$
Meanwhile, $\Delta_{p}u$ has the following expression (cf. [38, 41]),
$\Delta_{p}u=$ $f^{\frac{p}{2}-1}\left((p-1)u_{11}+\sum_{i=2}^{n}u_{ii}\right).$
Substituting the above equality into equation (3.2), we obtain:
(3.12) $(p-1)u_{11}+\sum_{i=2}^{n}u_{ii}=f+be^{cu}f^{1-\frac{p}{2}}.$
By Cauchy inequality, we arrive at
(3.13) $|\nabla\nabla u|^{2}\geq$ $\sum_{i=1}^{n}u_{1i}^{2}+\sum_{i=2}u_{ii}^{2}\geq\frac{|\nabla f|^{2}}{4f}+\frac{1}{n-1}\left(\sum_{i=2}u_{ii}\right)^{2}.$
It follows from (3.2) that
$\langle\nabla\Delta_{p}u,\nabla u\rangle=pf^{\frac{p}{2}}u_{11}+bce^{cu}f.$
Substituting (3.11), (3.13) and the above equality into (3.5) yields
(3.14) $\begin{split}\frac{f^{2-\alpha-\frac{p}{2}}}{2\alpha}{\mathscr{L}}\left(f^{\alpha}\right)\geq&\frac{1}{2}\left(\alpha+\frac{p-3}{2}\right)\frac{|\nabla f|^{2}}{f}+\frac{1}{n-1}\left(\sum_{i=2}u_{ii}\right)^{2}+\mathrm{Ric}(\nabla u,\nabla u)\\ &+2(p-2)(\alpha-1)u_{11}^{2}+f^{1-\frac{p}{2}}\left(pf^{\frac{p}{2}}u_{11}+bce^{cu}f\right).\end{split}$
Furthermore, by the facts that
$\frac{|\nabla f|^{2}}{f}\geq 4u_{11}^{2},\ \ \ \text{and}\ \ \ \alpha+\frac{p-3}{2}>0,$
we can infer from (3.14) that
(3.15) $\begin{split}\frac{f^{2-\alpha-\frac{p}{2}}}{2\alpha}{\mathscr{L}}\left(f^{\alpha}\right)\geq&2\left(\alpha+\frac{p-3}{2}\right)u_{11}^{2}+\frac{1}{n-1}\left(\sum_{i=2}^{n}u_{ii}\right)^{2}+\mathrm{Ric}(\nabla u,\nabla u)\\ &+2(p-2)(\alpha-1)u_{11}^{2}+f^{1-\frac{p}{2}}\left(pf^{\frac{p}{2}}u_{11}+bce^{cu}f\right).\end{split}$
By (3.12), we have
$\left(\sum_{i=2}^{n}u_{ii}\right)^{2}=$ $\left(f+be^{cu}f^{1-\frac{p}{2}}-(p-1)u_{11}\right)^{2}$
$=$ $f^{2}+\left(be^{cu}f^{1-\frac{p}{2}}-(p-1)u_{11}\right)^{2}+2be^{cu}f^{2-\frac{p}{2}}-2f(p-1)u_{11}.$
Substituting the above inequality into (3.15) yields
(3.16) $\begin{split}\frac{f^{2-\alpha-\frac{p}{2}}}{2\alpha}{\mathscr{L}}\left(f^{\alpha}\right)\geq&(p-1)(2\alpha-1)u_{11}^{2}-\mathrm{Ric}_{-}f+\left(p-\frac{2(p-1)}{n-1}\right)fu_{11}+\frac{f^{2}}{n-1}\\ &+b\left(c+\frac{2}{n-1}\right)e^{cu}f^{2-\frac{p}{2}}+\frac{1}{n-1}\left(be^{cu}f^{1-\frac{p}{2}}-(p-1)u_{11}\right)^{2}.\end{split}$
Now, denote
$a_{1}=\left|p-\frac{2(p-1)}{n-1}\right|.$
It follows from (3.11) that
$2\left(p-\frac{2(p-1)}{n-1}\right)fu_{11}\geq-a_{1}f^{\frac{1}{2}}|\nabla f|.$
Hence,
(3.17) $\begin{split}\frac{f^{2-\alpha-\frac{p}{2}}}{2\alpha}{\mathscr{L}}(f^{\alpha})\geq&(p-1)(2\alpha-1)u_{11}^{2}-\mathrm{Ric}_{-}f-\frac{a_{1}}{2}f^{\frac{1}{2}}|\nabla f|+\frac{f^{2}}{n-1}\\ &+b\left(c+\frac{2}{n-1}\right)e^{cu}f^{2-\frac{p}{2}}+\frac{1}{n-1}\left(be^{cu}f^{1-\frac{p}{2}}-(p-1)u_{11}\right)^{2}.\end{split}$
Case $1$: the constants $a$, $p$ and $q$ satisfy
$a\left(\frac{n+1}{n-1}-\frac{q}{p-1}\right)\geq 0.$
For this case we have
$be^{cu}f\left(c+\frac{2}{n-1}\right)=a(p-1)^{p-1}e^{cu}f\left(\frac{n+1}{n-1}-\frac{q}{p-1}\right)\geq 0.$
Since $\alpha\geq 1$, by discarding some non-negative terms in (3.17), we obtain
${\mathscr{L}}(f^{\alpha})\geq$ $2\alpha f^{\alpha+\frac{p}{2}-2}\left(\frac{f^{2}}{n-1}-\mathrm{Ric}_{-}f-\frac{a_{1}}{2}f^{\frac{1}{2}}|\nabla f|\right),$
which is just the inequality in the first case of Lemma 3.3.
By expanding the last term of the right hand side of (3.17), we obtain
(3.18) $\begin{split}\frac{f^{2-\alpha-\frac{p}{2}}}{2\alpha}{\mathscr{L}}(f^{\alpha})\geq&(p-1)\left(2\alpha-1+\frac{p-1}{n-1}\right)u_{11}^{2}-\mathrm{Ric}_{-}f\\ &+b\left(\frac{n+1}{n-1}-\frac{q}{p-1}\right)e^{cu}f^{2-\frac{p}{2}}-\frac{a_{1}}{2}f^{\frac{1}{2}}|\nabla f|+\frac{f^{2}}{n-1}\\ &+\frac{1}{n-1}\left(b^{2}e^{2cu}f^{2-p}-2(p-1)be^{cu}f^{1-\frac{p}{2}}u_{11}\right).\end{split}$
Case $2$ : the constants $a$, $p$ and $q$ satisfy
$p-1<q<\frac{n+3}{n-1}(p-1).$
In the present situation, the condition is equivalent to
$\frac{1}{n-1}-\frac{n-1}{4}\left(\frac{n+1}{n-1}-\frac{q}{p-1}\right)^{2}>0.$
This implies
$\lim_{\alpha\to\infty}\frac{1}{n-1}-\left(\frac{n+1}{n-1}-\frac{q}{p-1}\right)^{2}\frac{(2\alpha-1)(n-1)+p-1}{4(2\alpha-1)}>0.$
By the monotonicity of the left hand side of the above with repect to $\alpha$, it’s easy to see that if we denote
$\alpha_{0}(n,p,q)=\frac{\frac{4}{n-1}+(p-n)\left(\frac{n+1}{n-1}-\frac{q}{p-1}\right)^{2}}{2\left(\frac{4}{n-1}-(n-1)\left(\frac{n+1}{n-1}-\frac{q}{p-1}\right)^{2}\right)},$
then $\beta_{n,p,q,\alpha}>0$ when $\alpha>\alpha_{0}$.
On the other hand, by using the inequality $\lambda^{2}-2\lambda\mu\geq-\mu^{2}$ we have
(3.19) $\begin{split}&(p-1)\left(2\alpha-1+\frac{p-1}{n-1}\right)u_{11}^{2}-2\frac{(p-1)}{n-1}be^{cu}f^{1-\frac{p}{2}}u_{11}\\ \geq&-\frac{(p-1)b^{2}e^{2cu}f^{2-p}}{((2\alpha-1)(n-1)+p-1)(n-1)}.\end{split}$
Combining (3.18) and (3.19) yields
(3.20) $\begin{split}\frac{f^{2-\alpha-\frac{p}{2}}}{2\alpha}{\mathscr{L}}(f^{\alpha})\geq&\frac{(2\alpha-1)b^{2}e^{2cu}f^{2-p}}{(2\alpha-1)(n-1)+p-1}-\mathrm{Ric}_{-}f-\frac{a_{1}}{2}f^{\frac{1}{2}}|\nabla f|\\ &+b\left(\frac{n+1}{n-1}-\frac{q}{p-1}\right)e^{cu}f^{2-\frac{p}{2}}+\frac{f^{2}}{n-1}.\end{split}$
Applying the relation $\lambda^{2}+2\lambda\mu\geq-\mu^{2}$ again, we have
(3.21) $\begin{split}&\frac{(2\alpha-1)b^{2}e^{2cu}f^{2-p}}{(2\alpha-1)(n-1)+p-1}+b\left(\frac{n+1}{n-1}-\frac{q}{p-1}\right)e^{cu}f^{2-\frac{p}{2}}\\ \geq&-\left(\frac{n+1}{n-1}-\frac{q}{p-1}\right)^{2}\frac{(2\alpha-1)(n-1)+p-1}{4(2\alpha-1)}f^{2}.\end{split}$
Substituting (3.21) into (3.20), we arrive at
$\frac{f^{2-\alpha-\frac{p}{2}}}{2\alpha}{\mathscr{L}}(f^{\alpha})\geq$ $\left(\frac{1}{n-1}-\left(\frac{n+1}{n-1}-\frac{q}{p-1}\right)^{2}\frac{(2\alpha-1)(n-1)+p-1}{4(2\alpha-1)}\right)f^{2}$
$-\mathrm{Ric}_{-}f-\frac{a_{1}}{2}f^{\frac{1}{2}}|\nabla f|.$
Hence,
${\mathscr{L}}(f^{\alpha})\geq$ $2\beta_{n,p,q,\alpha}\alpha f^{\alpha+\frac{p}{2}}-2\alpha\mathrm{Ric}_{-}f^{\alpha+\frac{p}{2}-1}-a_{1}\alpha f^{\alpha+\frac{p}{2}-\frac{3}{2}}|\nabla f|,$
where $\beta_{n,p,q,\alpha}>0$ is defined in Lemma 3.3. Thus, we complete the proof of this lemma. ∎
3.C. Approximation procedure and key integral inequality
Now, we are going to establish a key integral inequality of $f=|\nabla u|^{2}$.
Lemma 3.4.
Let $\Omega=B_{R}(o)\subset M$ be a geodesic ball. Define $\alpha$ and $\beta_{n,p,q,\alpha}$ as in Lemma 3.3. Denote
$\theta(\alpha,t,p)=\alpha+t+\frac{p}{2}-1.$
Then, for
(3.22) $t\in\left(\frac{a_{1}^{2}}{a_{2}\beta_{n,p,q,\alpha}},\,\,\infty\right),$
the following inequality holds true
(3.23) $\frac{\beta_{n,p,q,\alpha}}{2}\int_{\Omega}f^{\theta+1}\eta^{2}+\frac{a_{2}t}{\theta^{2}}\mathbb{S}_{\chi}(M)\left\|f^{\theta}\eta^{2}\right\|_{L^{\chi}}\leq\int_{\Omega}\mathrm{Ric}_{-}f^{\theta}\eta^{2}+\gamma(\alpha,p,t)\int_{\Omega}f^{\theta}|\nabla\eta|^{2},$
where
$a_{1}=\left|p-\frac{2(p-1)}{n-1}\right|,\ \ \ a_{2}=\min\{1,\,p-1\},\ \ \ \text{and}\ \ \ \gamma(\alpha,p,t)=\frac{2(p+1)^{2}}{a_{2}t}+\frac{a_{2}t}{\theta^{2}}.$
Proof.
Since Lemma 3.3 only holds pointwisely on $\{x\in M:f(x)>0\}$. In order to obtain this integral estimate, we need to perform an approximation procedure as that in [65, 38]. Now, let $\eta\in C^{\infty}_{0}(\Omega,{\mathbb{R}})$ be a non-negative and smooth function on $\Omega$ with compact support. Denote $f_{\epsilon}=(f-\epsilon)^{+}$. By multiplying $\psi=f_{\epsilon}^{t}\eta^{2}$ on both side of (3.10)(where $t>1$ is to be determined later), we have
$-\int_{\Omega}\langle f^{p/2-1}\nabla f^{\alpha}+(p-2)f^{p/2-2}\langle\nabla f^{\alpha},\nabla u\rangle\nabla u,\,\nabla\psi\rangle$
$\geq$ $2\beta_{n,p,q,\alpha}\alpha\int_{\Omega}f^{\alpha+\frac{p}{2}}f^{t}_{\epsilon}\eta^{2}-2\alpha\int_{\Omega}\mathrm{Ric}_{-}f^{\alpha+\frac{p}{2}-1}f^{t}_{\epsilon}\eta^{2}-a_{1}\alpha\int_{\Omega}f^{\alpha+\frac{p-3}{2}}f^{t}_{\epsilon}|\nabla f|\eta^{2}.$
Hence,
(3.24) $\begin{split}&-\int_{\Omega}(\alpha tf^{\alpha+\frac{p}{2}-2}f^{t-1}_{\epsilon}|\nabla f|^{2}\eta^{2}+t\alpha(p-2)f^{\alpha+\frac{p}{2}-3}f^{t-1}_{\epsilon}\langle\nabla f,\nabla u\rangle^{2}\eta^{2})\\ &-\int_{\Omega}(2\eta\alpha f^{\alpha+\frac{p}{2}-2}f^{t}_{\epsilon}\langle\nabla f,\nabla\eta\rangle+2\alpha\eta(p-2)f^{\alpha+\frac{p}{2}-3}f^{t}_{\epsilon}\langle\nabla f,\nabla u\rangle\langle\nabla u,\nabla\eta\rangle)\\ \geq&2\beta_{n,p,q,\alpha}\alpha\int_{\Omega}f^{\alpha+\frac{p}{2}}f^{t}_{\epsilon}\eta^{2}-2\alpha\int_{\Omega}\mathrm{Ric}_{-}f^{\alpha+\frac{p}{2}-1}f^{t}_{\epsilon}\eta^{2}-a_{1}\alpha\int_{\Omega}f^{\alpha+\frac{p-3}{2}}f^{t}_{\epsilon}|\nabla f|\eta^{2}.\end{split}$
Notice that
(3.25) $f^{t-1}_{\epsilon}|\nabla f|^{2}+(p-2)f^{t-1}_{\epsilon}f^{-1}\langle\nabla f,\nabla u\rangle^{2}\geq a_{2}f^{t-1}_{\epsilon}|\nabla f|^{2},$
where $a_{2}=\min\{1,\,p-1\}$ and
(3.26) $f^{t}_{\epsilon}\langle\nabla f,\nabla\eta\rangle+(p-2)f^{t}_{\epsilon}f^{-1}\langle\nabla f,\nabla u\rangle\langle\nabla u,\nabla\eta\rangle\geq-(p+1)f^{t}_{\epsilon}|\nabla f||\nabla\eta|.$
Denote
$\theta=\alpha+t+\frac{p}{2}-1.$
Substituting (3.25) and (3.26) into (3.24), and then letting $\epsilon\to 0$, we arrive at
(3.27) $\begin{split}&2\beta_{n,p,q,\alpha}\int_{\Omega}f^{\theta+1}\eta^{2}+a_{2}t\int_{\Omega}f^{\theta-2}|\nabla f|^{2}\eta^{2}\\ \leq&2\int_{\Omega}\mathrm{Ric}_{-}f^{\theta}\eta^{2}+a_{1}\int_{\Omega}f^{\theta-\frac{1}{2}}|\nabla f|\eta^{2}+2(p+1)\int_{\Omega}f^{\theta-1}|\nabla f||\nabla\eta|\eta.\end{split}$
Since $u\in W^{2,2}_{loc}(\Omega\setminus\Omega_{cr})\cap C^{1,\beta}(\Omega)$ and the measure of critical set $\Omega_{cr}$ is zero by a very recent result [3, Corollary 1.6], we have $f\in C^{\beta}(\Omega)$ and $|\nabla f|\in L^{2}_{loc}$, and hence the integrals in the above make sense.
By Cauchy-inequality, we have
(3.28) $a_{1}f^{\theta-\frac{1}{2}}|\nabla f|\eta^{2}\leq\frac{a_{2}t}{4}f^{\theta-2}|\nabla f|^{2}\eta^{2}+\frac{a_{1}^{2}}{a_{2}t}f^{\theta+1}\eta^{2},$
and
(3.29) $2(p+1)f^{\theta-1}|\nabla f||\nabla\eta|\eta\leq\frac{a_{2}t}{4}f^{\theta-2}|\nabla f|^{2}\eta^{2}+\frac{4(p+1)^{2}}{a_{2}t}f^{\theta}|\nabla\eta|^{2}.$
Combining the fact
$t\in\left(\frac{a_{1}^{2}}{a_{2}\beta_{n,p,q,\alpha}},\,\,\infty\right),$
we conclude by substituting (3.28) and (3.29) into (3.27) that
(3.30) $\beta_{n,p,q,\alpha}\int_{\Omega}f^{\theta+1}\eta^{2}+\frac{a_{2}t}{2}\int_{\Omega}f^{\theta-2}|\nabla f|^{2}\eta^{2}\leq 2\int_{\Omega}\mathrm{Ric}_{-}f^{\theta}\eta^{2}+\frac{4(p+1)^{2}}{a_{2}t}\int_{\Omega}f^{\theta}|\nabla\eta|^{2}.$
On the other hand,
(3.31) $\begin{split}\frac{1}{2}\left|\nabla\left(f^{\frac{\theta}{2}}\eta\right)\right|^{2}\leq&\left|\nabla f^{\frac{\theta}{2}}\right|^{2}\eta^{2}+f^{\theta}|\nabla\eta|^{2}\\ =&\frac{\theta^{2}}{4}f^{\theta-2}|\nabla f|^{2}\eta^{2}+f^{\theta}|\nabla\eta|^{2}.\end{split}$
Substituting (3.31) into (3.30) yields
(3.32) $\beta_{n,p,q,\alpha}\int_{\Omega}f^{\theta+1}\eta^{2}+\frac{2a_{2}t}{\theta^{2}}\int_{\Omega}\left|\nabla\left(f^{\frac{\theta}{2}}\eta\right)\right|^{2}$
$\leq$ $2\int_{\Omega}\mathrm{Ric}_{-}f^{\theta}\eta^{2}+\left(\frac{4(p+1)^{2}}{a_{2}t}+\frac{2a_{2}t}{\theta^{2}}\right)\int_{\Omega}f^{\theta}|\nabla\eta|^{2}.$
By Sobolev inequality, there holds
$\mathbb{S}_{\chi}(M)\left\|f^{\frac{\theta}{2}}\eta\right\|_{L^{2\chi}(\Omega)}^{2}\leq\int_{\Omega}\left|\nabla\left(f^{\frac{\theta}{2}}\eta\right)\right|^{2}.$
Hence,
(3.33) $\begin{split}&\frac{\beta_{n,p,q,\alpha}}{2}\int_{\Omega}f^{\theta+1}\eta^{2}+\frac{a_{2}t}{\theta^{2}}\mathbb{S}_{\chi}(M)\left\|f^{\theta}\eta^{2}\right\|_{L^{\chi}}\\ \leq&\int_{\Omega}\mathrm{Ric}_{-}f^{\theta}\eta^{2}+\left(\frac{2(p+1)^{2}}{a_{2}t}+\frac{a_{2}t}{\theta^{2}}\right)\int_{\Omega}f^{\theta}|\nabla\eta|^{2}.\end{split}$
We complete the proof. ∎
3.D. Local $L^{\theta\chi}$ bound of the gradient.
Next, we are ready to show the following $L^{\theta\chi}$ bound with
$\theta=\alpha+t+p/2-1$
of the gradient of positive solutions to equation (1.1).
Lemma 3.5.
Let $(M,g)$ be a complete manifold on which the $\chi$-type Sobolev inequality holds. Assume $u$ is a positive solution to equation (3.2) on the geodesic ball $B(o,R)\subset M$. Let $f=|\nabla u|^{2}$ and denote $\theta$ by
$\theta=\alpha+t+\frac{p}{2}-1,$
where $\alpha$ and $t$ satisfy the conditions in Lemma 3.3 and (3.22) respectively. Assume further that
$\beta(\alpha,p,t)\triangleq\frac{a_{2}t}{\theta^{2}}\mathbb{S}_{\chi}(M)-\left\|\mathrm{Ric}_{-}\right\|_{L^{\frac{\chi}{\chi-1}}}>0.$
Then there exists $a_{3}=a_{3}(n,\alpha,p,q,t)>0$ such that
(3.34) $\|f\|_{L^{\theta\chi}(B_{3R/4}(o))}\leq a_{3}\left(\frac{V}{R^{2(\theta+1)}}\right)^{\frac{1}{\theta}},$
where $V$ denotes the volume of geodesic ball $B_{R}(o)$.
Proof.
By Hölder’s inequality, we have
$\int_{\Omega}\mathrm{Ric}_{-}f^{\theta}\eta^{2}\leq\left\|\mathrm{Ric}_{-}\right\|_{L^{\frac{\chi}{\chi-1}}}\left\|f^{\theta}\eta^{2}\right\|_{L^{\chi}}.$
Substituting this into Lemma 3.4, we conclude that
(3.35) $\frac{\beta_{n,p,q,\alpha}}{2}\int_{B(o,R)}f^{\theta+1}\eta^{2}+\left(\frac{a_{2}t}{\theta^{2}}\mathbb{S}_{\chi}(M)-\left\|\mathrm{Ric}_{-}\right\|_{L^{\frac{\chi}{\chi-1}}}\right)\left\|f^{\theta}\eta^{2}\right\|_{L^{\chi}}$
$\leq$ $\gamma(\alpha,p,t)\int_{B(o,R)}f^{\theta}|\nabla\eta|^{2},$
where
$\gamma(\alpha,p,t)=\frac{2(p+1)^{2}}{a_{2}t}+\frac{a_{2}t}{\theta^{2}}.$
Now, choose $\eta_{1}\in C^{\infty}_{0}(B_{R}(o))$ such that
$\begin{cases}0\leq\eta_{1}\leq 1,\quad\eta_{1}\equiv 1\text{ in }B_{{3R}/{4}}(o);\\ |\nabla\eta_{1}|\leq\frac{C(n)}{R},\end{cases}$
and let $\eta=\eta_{1}^{\theta+1}$. Direct calculation shows that
$R^{2}|\nabla\eta|^{2}\leq C^{2}(n)\left(\theta+1\right)^{2}\eta^{\frac{2\theta}{\theta+1}}.$
By Hölder inequality and Young inequality, we have
(3.36) $\begin{split}\gamma(\alpha,p,t)\int_{B(o,R)}f^{\theta}|\nabla\eta|^{2}\leq&\frac{C^{2}(n)\left(\theta+1\right)^{2}\gamma(\alpha,p,t)}{R^{2}}\int_{B(o,R)}f^{\theta}\eta^{\frac{2\theta}{\theta+1}}\\ \leq&\frac{C^{2}(n)\left(\theta+1\right)^{2}\gamma(\alpha,p,t)}{R^{2}}\left(\int_{B(o,R)}f^{\theta+1}\eta^{2}\right)^{\frac{\theta}{\theta+1}}V^{\frac{1}{\theta+1}}\\ \leq&\frac{\beta_{n,p,q,\alpha}}{2}\left[\int_{B(o,R)}f^{\theta+1}\eta^{2}+\left(\frac{C^{2}(n)\left(\theta+1\right)^{2}\gamma(\alpha,p,t)}{\beta_{n,p,q,\alpha}R^{2}}\right)^{\theta+1}V\right].\end{split}$
Since $\eta\equiv 1$ in $B_{3R/4}$, there holds
(3.37) $\left\|f^{\theta}\right\|_{L^{\chi}(B_{3R/4}(o))}\leq\left\|f^{\theta}\eta^{2}\right\|_{L^{\chi}(B_{R}(o))}.$
Substituting (3.36) and (3.37) into (3.35) and keeping in mind
$\left(\frac{a_{2}t}{\theta^{2}}\mathbb{S}_{\chi}(M)-\left\|\mathrm{Ric}_{-}\right\|_{L^{\frac{\chi}{\chi-1}}}\right)>0,$
we arrive at
(3.38) $\left\|f^{\theta}\right\|_{L^{\chi}(B_{3R/4}(o))}\leq\frac{\beta_{n,p,q,\alpha}}{\beta(\alpha,p,t)}\left(\frac{C^{2}(n)\left(\theta+1\right)^{2}\gamma(\alpha,p,t)}{\beta_{n,p,q,\alpha}R^{2}}\right)^{\theta+1}V.$
Finally, we conclude that
$\|f\|_{L^{\theta\chi}(B_{3R/4}(o))}\leq\frac{\left(C^{2}(n)\left(\theta+1\right)^{2}\gamma(\alpha,p,t)\right)^{\frac{\theta+1}{\theta}}}{\beta_{n,p,q,\alpha}\beta(\alpha,p,t)^{\frac{1}{\theta}}}\left(\frac{V}{R^{2(\theta+1)}}\right)^{\frac{1}{\theta}}.$
Proof of Theorem 1.4:.
Firstly, fix some $\alpha>0$ such that $\alpha$ meets the conditions in Lemma 3.3. Then we can choose a large $t$ such that
$2\left(\alpha+t+\frac{p}{2}\right)>\frac{2\chi}{\chi-1}\quad\mbox{as}\,\,\chi\in\left(1,\,\frac{n}{n-2}\right)$
or
$2\left(\alpha+t+\frac{p}{2}\right)>\beta^{*}\geq n\quad\mbox{as}\,\,\chi=\frac{n}{n-2},$
and the condition (3.22) holds. Once these have been done, we see that there exists a positive constant $C(n,p,q,\beta^{*})$ depending on $n,p,q$ and $\beta^{*}$ such that, if
$\|\mathrm{Ric}_{-}\|_{L^{\frac{\chi}{\chi-1}}}<C(n,p,q,\beta^{*})\mathbb{S}_{\chi}(M),$
then relation (3.34) holds. Now, by letting $R$ tends to infinity, we conclude that $f=0$, i.e., $\nabla u=0$. Thus, $v$ is a positive constant. This contradicts to $v$ is a solution. We complete the proof. ∎
3.E. Local gradient estimate when $\mathrm{Ric}_{-}\in L^{\gamma}$ for some $\gamma>\chi/(\chi-1)$.
Theorem 3.6.
Let $(M,g)$ be a complete noncompact Riemannian manifold on which the $\chi$-type Sobolev inequality holds. Denote $\Lambda=\|\mathrm{Ric}_{-}\|_{L^{\gamma}(B_{1})}$ for some $\gamma>\chi/(\chi-1)$. Then, for any $r\leq 1$ when
$a>0\quad\&\quad q<\frac{n+3}{n-1}(p-1)\quad\text{or}\quad a<0\quad\&\quad q>p-1,$
the following local gradient estimate holds for positive solution $v$ to (1.1),
$\sup_{B_{r/2}}\frac{|\nabla v|^{2}}{v^{2}}\leq a_{5}\left(\frac{V}{r^{2\left(\frac{\chi}{\chi-1}+\theta\right)}}\right)^{\frac{1}{\theta}},$
where $V$ is the volume of $B_{r}$.
Proof.
Notice that, by Lemma 3.4, we already have
(3.39) $\frac{\beta_{n,p,q,\alpha}}{2}\int_{B_{r}}f^{\theta+1}\eta^{2}+\frac{a_{2}t}{\theta^{2}}\mathbb{S}_{\chi}(M)\left\|f^{\theta}\eta^{2}\right\|_{L^{\chi}}\leq\int_{B_{r}}\mathrm{Ric}_{-}f^{\theta}\eta^{2}+\gamma(\alpha,p,t)\int_{B_{r}}f^{\theta}|\nabla\eta|^{2}.$
By Holder’s inequality, we have
(3.40) $\int_{\Omega}\mathrm{Ric}_{-}f^{\theta}\eta^{2}\leq\Lambda\|f^{\frac{\theta}{2}}\eta\|_{L^{\frac{2\gamma}{\gamma-1}}}^{2}.$
Notice that
$\frac{2\gamma}{\gamma-1}\in(2,\,2\chi).$
By interpolation inequality, we have
$\left\|f^{\frac{\theta}{2}}\eta\right\|_{L^{\frac{2\gamma}{\gamma-1}}}\leq\varepsilon\left\|f^{\frac{\theta}{2}}\eta\right\|_{L^{2\chi}}+\varepsilon^{-\frac{\chi}{(\chi-1)\gamma-\chi}}\left\|f^{\frac{\theta}{2}}\eta\right\|_{L^{2}}.$
Hence
$\int_{\Omega}\mathrm{Ric}_{-}f^{\theta}\eta^{2}\leq 2\Lambda\varepsilon^{2}\left\|f^{\frac{\theta}{2}}\eta\right\|_{L^{2\chi}}^{2}+2\Lambda\varepsilon^{-\frac{2\chi}{(\chi-1)\gamma-\chi}}\left\|f^{\frac{\theta}{2}}\eta\right\|_{L^{2}}^{2}.$
Without loss of generality, we assume $\Lambda>0$. Now, Let
$\varepsilon=\varepsilon(\alpha,p,t)=\frac{1}{2\theta}\left(\frac{a_{2}t\mathbb{S}_{\chi}(M)}{\Lambda}\right)^{\frac{1}{2}}.$
Substituting the above into (3.39), we arrive at
(3.41) $\frac{\beta_{n,p,q,\alpha}}{2}\int_{B_{r}}f^{\theta+1}\eta^{2}+\frac{a_{2}t}{2\theta^{2}}\mathbb{S}_{\chi}(M)\left\|f^{\theta}\eta^{2}\right\|_{L^{\chi}}$
$\leq\,$ $\gamma(\alpha,p,t)\int_{B_{r}}f^{\theta}|\nabla\eta|^{2}+2\Lambda\varepsilon^{-\frac{2\chi}{(\chi-1)\gamma-\chi}}\int_{B_{r}}f^{\theta}\eta^{2}.$
By almost the same arguments as the proof of Lemma 3.5, we conclude that when $r\leq 1$, there exists $a_{3}=a_{3}(n,\alpha,p,q,t,\Lambda)>0$ such that
(3.42) $\|f\|_{L^{\theta\chi}(B_{3r/4}(o))}\leq a_{3}\left(\frac{V}{r^{2(\theta+1)}}\right)^{\frac{1}{\theta}},$
where $V$ denotes the volume of geodesic ball $B_{r}(o)$.
Now, we fix $\alpha=\alpha_{0}$ such that $\alpha_{0}$ satisfies the conditions in Lemma 3.3. Once $\alpha$ has been fixed, $\theta$ can be regarded as a function with respect to $t$. That is to say, when $\alpha=\alpha_{0}$ is fixed,
$\theta=\theta(t)=\alpha_{0}+t+\frac{p}{2}-1.$
On the other hand, when $r\leq 1$, the test function $\eta$ constructed below satisfies
$1=\|\eta\|_{L^{\infty}}\leq\|\nabla\eta\|_{L^{\infty}}.$
Hence, (3.41) can be rewritten as
(3.43) $\frac{\beta_{n,p,q,\alpha}}{2}\int_{\Omega}f^{\theta+1}\eta^{2}+\frac{a_{2}t}{2\theta^{2}}\mathbb{S}_{\chi}(M)\left\|f^{\theta}\eta^{2}\right\|_{L^{\chi}}\leq\left(\gamma(\alpha,p,t)+2\Lambda\varepsilon^{-\frac{2\chi}{(\chi-1)\gamma-\chi}}\right)|\nabla\eta|_{L^{\infty}}^{2}\int_{\Omega}f^{\theta}.$
Now, denote by
$\zeta(t)=\frac{2\theta^{2}\left(\gamma(\alpha_{0},p,t)+2\Lambda\varepsilon^{-\frac{2\chi}{(\chi-1)\gamma-\chi}}\right)}{a_{2}t\mathcal{S}_{\chi}(M)}.$
It follows from (3.43) that
(3.44) $\left\|f^{\theta}\eta^{2}\right\|_{L^{\chi}}\leq\zeta(t)|\nabla\eta|_{L^{\infty}}^{2}\int_{\Omega}f^{\theta}.$
Let
$\Omega_{k}=B_{r_{k}}(o)\quad\text{with}\quad r_{k}=\frac{r}{2}+\frac{r}{4^{k}},$
and choose $\eta_{k}\in C^{\infty}(\Omega_{k})$ satisfying
$\begin{cases}0\leq\eta_{k}\leq 1,\quad\eta_{k}\equiv 1\text{ in }B_{r_{k+1}}(o);\\ |\nabla\eta_{k}|\leq\frac{C4^{k}}{r}.\end{cases}$
Now, for any $t_{0}$ satisfying the condition in Lemma 3.4, we denote
$\theta_{0}=\alpha_{0}+t_{0}+\frac{p}{2}-1.$
Moreover, we let $\beta_{k}=\theta_{0}\chi^{k}$ and $t=t_{k}$ such that
$t_{k}+\frac{p}{2}+\alpha_{0}-1=\beta_{k}.$
By substituting $\theta=\theta_{k}$ and $\eta$ by $\eta_{k}$ in (3.44), we arrive at
$\|f\|_{L^{\beta_{k+1}}(\Omega_{k+1})}\leq\zeta(t_{k})^{\frac{1}{\beta_{k}}}16^{\frac{k}{\beta_{k}}}r^{-\frac{2}{\beta_{k}}}\|f\|_{L^{\beta_{k}}(\Omega_{k})}.$
Hence,
(3.45) $\|f\|_{L^{\beta_{k+1}}(\Omega_{k+1})}\leq\prod_{i=1}^{k}\zeta(t_{i})^{\frac{1}{\beta_{i}}}16^{\sum_{i=1}^{k}\frac{i}{\beta_{i}}}r^{-\sum_{i=1}^{k}\frac{2}{\beta_{i}}}\|f\|_{L^{\beta_{1}}(\Omega_{k})}.$
Notice that
$\zeta(t)\leq c(p,\alpha_{0},\Lambda,\mathbb{S}_{\chi}(M))t^{\frac{(\chi-1)\gamma}{(\chi-1)\gamma-\chi}}.$
Straightforward calculation shows that
$\prod_{i=1}^{\infty}\zeta(t_{i})^{\frac{1}{\beta_{i}}}<\infty,\quad\quad\sum_{k=1}^{\infty}\frac{1}{\beta_{k}}=\frac{1}{\theta_{0}(\chi-1)},\quad\quad\sum_{k=1}^{\infty}\frac{k}{\beta_{k}}<\infty.$
By letting $k\rightarrow\infty$ in (3.45), we arrive at
$\|f\|_{L^{\infty}(B_{r/2}(o))}\leq$ $a_{4}r^{-\frac{2}{\theta_{0}(\chi-1)}}\|f\|_{L^{\beta}(B_{3r/4}(o))}.$
By substituting (3.42) into the above, we finally arrive at
$\|f\|_{L^{\infty}(B_{r/2}(o))}\leq a_{5}(\alpha_{0},t_{0},p,q,\Lambda,\mathbb{S}_{\chi}(M))\left(\frac{V}{r^{2\left(\frac{\chi}{\chi-1}+\theta_{0}\right)}}\right)^{\frac{1}{\theta_{0}}}.$
Thus, we complete the proof. ∎
Proof of Theorem 1.9:.
Since $\chi\leq n/(n-2)$, it is easy to see that $\chi/(\chi-1)\geq n/2$. Hence $\gamma>n/2$. By the relative volume comparison theorem under the integral bounded Ricci curvature due to Peterson and Wei [54], there holds
(3.46) ${\rm{Vol}}(B_{r})\leq\omega_{n}r^{n}+C(n,\gamma)\|\mathrm{Ric}_{-}\|^{\gamma}_{L^{\gamma}}r^{2\gamma}.$
Hence, by substituting the above into Theorem 3.6, then letting $r=1$, we conclude the conclusion. ∎
4. Laplace Case: Proof of Theorem 1.5
In the previous section (see Theorem 1.4), we have shown that the conclusion of Theorem 1.5 holds for $q<\frac{n+3}{n-1}$. In this section, we shall prove the remaining case of
$q\in\left[\frac{n+3}{n-1},\penalty 10000\ \frac{n+2}{(n-2)_{+}}\right).$
Throughout this section, we assume $v$ be a positive solution of (1.2) on $B(o,R)\subset M$. First, we recall some auxiliary functions and point-wise estimates developed by Lu in [47]. As categorized in [47], we need to consider dimensions greater than or equal to $4$ and less than $4$ respectively.
4.A. Case 1: $n\geq 4\quad\&\quad q\in\left[\frac{n+3}{n-1},\penalty 10000\ \frac{n+2}{n-2}\right)$
4.A.1. First auxiliary function $F$ and estimation of the leading coefficients.
For $\theta\neq 0$, let $\omega=v^{-\theta}$. A straightword calculation shows that
(4.1) $\Delta\omega=\left(1+\frac{1}{\theta}\right)\frac{|\nabla\omega|^{2}}{\omega}+\theta\omega v^{q-1}.$
For undetermined real numbers $\varepsilon>0$ and $d>0$, define the first type auxiliary function:
(4.2) $F=(v+\varepsilon)^{-\theta}\left(\frac{|\nabla\omega|^{2}}{\omega^{2}}+dv^{q-1}\right).$
Lemma 4.1.
(Lemma 2.1($\mathrm${Lemma\,2.1} and Lemma 6.1 in [47])) There holds:
(4.3) $\begin{split}(v+\varepsilon)^{\theta}\Delta F=\,&2\omega^{-2}\left|\nabla^{2}\omega-\frac{\Delta\omega}{n}g\right|^{2}+2\omega^{-2}\mathrm{Ric}(\nabla\omega,\nabla\omega)\\ &+2\left(\frac{1}{\theta}-\frac{\varepsilon}{v+\varepsilon}\right)(v+\varepsilon)^{\theta}\langle\nabla F,\nabla\ln\omega\rangle\\ &+U\frac{|\nabla\omega|^{4}}{\omega^{4}}+V\frac{|\nabla\omega|^{2}}{\omega^{2}}v^{q-1}+Wv^{2(q-1)},\end{split}$
where
$U=$ $\frac{2}{n}\left(1+\frac{1}{\theta}\right)^{2}+\left(\frac{1}{\theta}-1\right)\frac{v^{2}}{(v+\varepsilon)^{2}}+2\left(1-\frac{1}{\theta}\right)\frac{v}{v+\varepsilon}-2,$
$V=$ $\frac{4}{n}(1+\theta)+2(1-q)+\frac{d(q-1)}{\theta^{2}}(q-2\theta)+\frac{v}{v+\varepsilon}\left\{\theta-d\left(\frac{1}{\theta}-1\right)\left(1+\frac{\varepsilon}{v+\varepsilon}\right)\right\},$
$W=$ $\frac{2\theta^{2}}{n}+d\left(\frac{\theta v}{v+\varepsilon}+1-q\right).$
Moreover, for $n\geq 4$ and $q\in\left[\frac{n+3}{n-1},\penalty 10000\ \frac{n+2}{n-2}\right)$, there exist $\theta=\theta(n,q)\in\left(0,\frac{2}{n-2}\right)$, $d=d(n,q)>0$, $L=L(n,q)>0$ and $\widetilde{M}=\widetilde{M}(n,q)>0$ such that for any $\varepsilon>0$,
$U$ $\geq U_{0}>0$
$V$ $\geq V_{0}-\widetilde{M}\overline{\chi}_{\left\{x\in B(o,R):\,v(x)<L\varepsilon\right\}},$
$W$ $\geq W_{0}-\widetilde{M}\overline{\chi}_{\left\{x\in B(o,R):\,v(x)<L\varepsilon\right\}},$
where $U_{0}$, $V_{0}$, $W_{0}$ are positive constants depending only on $n$ and $q$, and $\overline{\chi}$ denotes the characteristic function.
Next, we shall provide a precise estimate for $\Delta F$.
Lemma 4.2.
Let $n\geq 4$ and $q\in\left[\frac{n+3}{n-1},\penalty 10000\ \frac{n+2}{n-2}\right)$. Then there exist $\theta=\theta(n,q)\in\left(0,\frac{2}{n-2}\right)$, $d=d(n,q)>0$, $C_{0}=C_{0}(n,q)>0$ and $M_{0}=M_{0}(n,q)>0$, such that the following holds point-wisely in $B(o,R)$:
(4.4) $\Delta F\geq-2\mathrm{Ric}_{-}F-\frac{2}{\theta}|\nabla F||\nabla\ln\omega|+C_{0}(v+\varepsilon)^{\theta}F^{2}-M_{0}\varepsilon^{q-1}F.$
Proof.
By Lemma 4.1, we know that there exist constants
$\theta=\theta(n,q)\in\left(0,\frac{2}{n-2}\right),\quad d=d(n,q)>0,\quad L=L(n,q)>0\quad\mbox{and}\quad\widetilde{M}=\widetilde{M}(n,q)>0$
such that for any $\varepsilon>0$,
(4.5) $\begin{split}(v+\varepsilon)^{\theta}\Delta F\geq\,&2\omega^{-2}\mathrm{Ric}(\nabla\omega,\nabla\omega)+2\left(\frac{1}{\theta}-\frac{\varepsilon}{v+\varepsilon}\right)(v+\varepsilon)^{\theta}\langle\nabla F,\nabla\ln\omega\rangle\\ &+U_{0}\frac{|\nabla\omega|^{4}}{\omega^{4}}+V_{0}\frac{|\nabla\omega|^{2}}{\omega^{2}}v^{q-1}+W_{0}v^{2(q-1)}\\ &-\widetilde{M}v^{q-1}\left(\frac{|\nabla\omega|^{2}}{\omega^{2}}+v^{q-1}\right)\overline{\chi}_{\left\{x\in B(o,R):v(x)<L\varepsilon\right\}},\end{split}$
where $U_{0}$, $V_{0}$, $W_{0}$ are positive constants that depend only on $n$ and $q$. Notice that
(4.6) $2\left(\frac{1}{\theta}-1+\frac{v}{v+\varepsilon}\right)(v+\varepsilon)^{\theta}\langle\nabla F,\nabla\ln\omega\rangle\geq-\frac{2}{\theta}(v+\varepsilon)^{\theta}|\nabla F||\nabla\ln\omega|,$
(4.7) $\begin{split}U_{0}\frac{|\nabla\omega|^{4}}{\omega^{4}}+V_{0}\frac{|\nabla\omega|^{2}}{\omega^{2}}v^{q-1}+W_{0}v^{2(q-1)}\geq&\min\left\{U_{0},\frac{V_{0}}{2},W_{0}\right\}\left(\frac{|\nabla\omega|^{2}}{\omega^{2}}+v^{q-1}\right)^{2}\\ \geq&C_{0}(n,q)\left(\frac{|\nabla\omega|^{2}}{\omega^{2}}+dv^{q-1}\right)^{2}\\ =&C_{0}(n,q)(v+\varepsilon)^{2\theta}F^{2},\end{split}$
and
(4.8) $\begin{split}-\widetilde{M}v^{q-1}\left(\frac{|\nabla\omega|^{2}}{\omega^{2}}+v^{q-1}\right)\chi_{\left\{x\in B(o,R):v(x)<L\varepsilon\right\}}\geq&-\widetilde{M}(L\varepsilon)^{q-1}\left(\frac{|\nabla\omega|^{2}}{\omega^{2}}+v^{q-1}\right)\\ \geq&-M_{0}(n,q)\varepsilon^{q-1}\left(\frac{|\nabla\omega|^{2}}{\omega^{2}}+dv^{q-1}\right)\\ =&-M_{0}(n,q)\varepsilon^{q-1}(v+\varepsilon)^{\theta}F,\end{split}$
where $C_{0}(n,q)$ and $M_{0}(n,q)$ are positive numbers and depend only on $n$ and $q$.
Now, substituting (4.6), (4.7) and (4.8) into (4.5) and dividing the both sides by $(v+\varepsilon)^{\theta}$ yields
$\Delta F\geq\,$ $2(v+\varepsilon)^{-\theta}\omega^{-2}\mathrm{Ric}(\nabla\omega,\nabla\omega)-2\left(\frac{1}{\theta}+1\right)|\nabla F||\nabla\ln\omega|$
(4.9) $+C_{0}(v+\varepsilon)^{\theta}F^{2}-M_{0}\varepsilon^{q-1}F.$
On the other hand, from (4.2) we obtain
$2(v+\varepsilon)^{-\theta}\omega^{-2}\mathrm{Ric}(\nabla\omega,\nabla\omega)\geq$ $-2(v+\varepsilon)^{-\theta}\mathrm{Ric_{-}}\frac{|\nabla\omega|^{2}}{\omega^{2}}\geq-2\mathrm{Ric_{-}}F.$
Substituting the above inequality into (4.A.1), we finish the proof of Lemma 4.2. ∎
4.A.2. Integral estimate
Now, we are going to establish a key integral inequality of $F$.
Lemma 4.3.
Let $(M,g)$ be a complete manifold on which the $\chi$-type Sobolev inequality holds. Let $n\geq 4$, $q\in\left[\frac{n+3}{n-1},\penalty 10000\ \frac{n+2}{n-2}\right)$ and $\Omega=B(o,R)$. Define $\theta$, $d$, $C_{0}$ and $M_{0}$ as in Lemma 4.2. Then, for
(4.10) $t\in\left(\max\left\{\frac{8}{C_{0}\theta^{2}},1\right\},\penalty 10000\ +\infty\right),$
the following holds
(4.11) $\begin{split}&C_{0}\varepsilon^{\theta}\int_{\Omega}F^{t+2}\eta^{2}+\frac{\mathbb{S}_{\chi}(M)}{2t}\left\|F^{t+1}\eta^{2}\right\|_{L^{\chi}(\Omega)}\\ \leq&4\int_{\Omega}\mathrm{Ric}_{-}F^{t+1}\eta^{2}+2M_{0}\varepsilon^{q-1}\int_{\Omega}F^{t+1}\eta^{2}+\frac{12}{t}\int_{\Omega}F^{t+1}|\nabla\eta|^{2},\end{split}$
where $\eta\geq 0$ and $\eta\in C_{0}^{\infty}(\Omega)$.
Proof.
Let $\eta\in C_{0}^{\infty}(\Omega)$ be a nonnegative function. By multiplying $F^{t}\eta^{2}$ on the both sides of (4.4) ($t>1$ will be determined later) and integration by parts, we arrive at
(4.12) $\begin{split}&2\int_{\Omega}\mathrm{Ric}_{-}F^{t+1}\eta^{2}+\frac{2}{\theta}\int_{\Omega}F^{t}|\nabla F||\nabla\ln\omega|\eta^{2}+M_{0}\varepsilon^{q-1}\int_{\Omega}F^{t+1}\eta^{2}-2\int_{\Omega}F^{t}\langle\nabla F,\nabla\eta\rangle\eta\\ \geq&t\int_{\Omega}F^{t-1}|\nabla F|^{2}\eta^{2}+C_{0}\int_{\Omega}(v+\varepsilon)^{\theta}F^{t+2}\eta^{2}.\end{split}$
Notice that
$\frac{2}{\theta}\int_{\Omega}F^{t}|\nabla F||\nabla\ln\omega|\eta^{2}\leq$ $\frac{t}{4}\int_{\Omega}F^{t-1}|\nabla F|^{2}\eta^{2}+\frac{4}{t\theta^{2}}\int_{\Omega}F^{t+1}|\nabla\ln\omega|^{2}\eta^{2}$
$\leq$ $\frac{t}{4}\int_{\Omega}F^{t-1}|\nabla F|^{2}\eta^{2}+\frac{4}{t\theta^{2}}\int_{\Omega}(v+\varepsilon)^{\theta}F^{t+2}\eta^{2}$
and
$-2\int_{\Omega}F^{t}\langle\nabla F,\nabla\eta\rangle\eta\leq\frac{t}{4}\int_{\Omega}F^{t-1}|\nabla F|^{2}\eta^{2}+\frac{4}{t}\int_{\Omega}F^{t+1}|\nabla\eta|^{2}.$
Substituting the above two inequalities into (4.12), we conclude
$\frac{t}{2}\int_{\Omega}F^{t-1}|\nabla F|^{2}\eta^{2}+\left(C_{0}-\frac{4}{t\theta^{2}}\right)\int_{\Omega}(v+\varepsilon)^{\theta}F^{t+2}\eta^{2}$
$\leq$ $2\int_{\Omega}\mathrm{Ric}_{-}F^{t+1}\eta^{2}+M_{0}\varepsilon^{q-1}\int_{\Omega}F^{t+1}\eta^{2}+\frac{4}{t}\int_{\Omega}F^{t+1}|\nabla\eta|^{2}.$
Let
$t\in\left(\max\left\{\frac{8}{C_{0}\theta^{2}},1\right\},\penalty 10000\ +\infty\right),$
then we have
(4.13) $\begin{split}&t\int_{\Omega}F^{t-1}|\nabla F|^{2}\eta^{2}+C_{0}\int_{\Omega}(v+\varepsilon)^{\theta}F^{t+2}\eta^{2}\\ \leq&4\int_{\Omega}\mathrm{Ric}_{-}F^{t+1}\eta^{2}+2M_{0}\varepsilon^{q-1}\int_{\Omega}F^{t+1}\eta^{2}+\frac{8}{t}\int_{\Omega}F^{t+1}|\nabla\eta|^{2}.\end{split}$
By $\chi$-type Sobolev inequality, there holds
$\mathbb{S}_{\chi}(M)\left\|F^{\frac{t+1}{2}}\eta\right\|_{L^{2\chi}(\Omega)}^{2}\leq\int_{\Omega}\left|\nabla\left(F^{\frac{t+1}{2}}\eta\right)\right|^{2}.$
Hence,
$\mathbb{S}_{\chi}(M)\left\|F^{t+1}\eta^{2}\right\|_{L^{\chi}(\Omega)}^{2}\leq\frac{(t+1)^{2}}{2}\int_{\Omega}F^{t-1}|\nabla F|^{2}\eta^{2}+2\int_{\Omega}F^{t+1}|\nabla\eta|^{2}.$
Substituting the above inequality into (4.13), then we obtain
$\mathbb{S}_{\chi}(M)\frac{2t}{(t+1)^{2}}\left\|F^{t+1}\eta^{2}\right\|_{L^{\chi}(\Omega)}^{2}+C_{0}\int_{\Omega}(v+\varepsilon)^{\theta}F^{t+2}\eta^{2}$
$\leq$ $4\int_{\Omega}\mathrm{Ric}_{-}F^{t+1}\eta^{2}+2M_{0}\varepsilon^{q-1}\int_{\Omega}F^{t+1}\eta^{2}+\left[\frac{8}{t}+\frac{4t}{(t+1)^{2}}\right]\int_{\Omega}F^{t+1}|\nabla\eta|^{2}.$
Therefore, we obtain
$\frac{\mathbb{S}_{\chi}(M)}{2t}\left\|F^{t+1}\eta^{2}\right\|_{L^{\chi}(\Omega)}^{2}+C_{0}\varepsilon^{\theta}\int_{\Omega}F^{t+2}\eta^{2}$
$\leq$ $4\int_{\Omega}\mathrm{Ric}_{-}F^{t+1}\eta^{2}+2M_{0}\varepsilon^{q-1}\int_{\Omega}F^{t+1}\eta^{2}+\frac{12}{t}\int_{\Omega}F^{t+1}|\nabla\eta|^{2}.$
Combining above, we finish the proof of Lemma 4.3.
Next, we shall provide the following $L^{(t+1)\chi}$ bound of $(v+1)^{-\theta}v^{q-1}$.
Lemma 4.4.
Let $(M,g)$ be a complete manifold on which the $\chi$-type Sobolev inequality holds. Assume $v$ is a positive solution to (1.2) on the geodesic ball $B(o,R)\subset M$. Furthermore, let $n\geq 4$, $q\in\left[\frac{n+3}{n-1},\penalty 10000\ \frac{n+2}{n-2}\right)$ and $\Omega=B(o,R)$. Define $t$ and $\theta$ as in Lemma 4.3. Assume further that
$H(t)=\frac{\mathbb{S}_{\chi}(M)}{2t}-4\left\|\mathrm{Ric_{-}}\right\|_{L^{\frac{\chi}{\chi-1}}}>0.$
Then there exists $C_{3}=C_{3}(n,q,t)>0$ such that
(4.14) $H(t)\left\{\int_{B\left(o,\frac{3}{4}R\right)}\left[(v+1)^{-\theta}v^{q-1}\right]^{(t+1)\chi}\right\}^{\frac{1}{\chi}}\leq C_{3}\left(\frac{1}{R^{2(t+2)}\varepsilon^{\theta(t+1)}}+\varepsilon^{(q-1)+(q-\theta-1)(t+1)}\right)V,$
where $V$ denotes the volume of geodesic ball $B(o,R)$ and $\varepsilon\in(0,1)$ is any positive number.
Proof.
By Hölder’s inequality, we have
$\int_{\Omega}\mathrm{Ric}_{-}F^{t+1}\eta^{2}\leq\left\|\mathrm{Ric_{-}}\right\|_{L^{\frac{\chi}{\chi-1}}}\left\|F^{t+1}\eta^{2}\right\|_{L^{\chi}(\Omega)}.$
Substituting this into (4.11), we conclude that
(4.15) $C_{0}\varepsilon^{\theta}\int_{\Omega}F^{t+2}\eta^{2}+H(t)\left\|F^{t+1}\eta^{2}\right\|_{L^{\chi}(\Omega)}\leq 2M_{0}\varepsilon^{q-1}\int_{\Omega}F^{t+1}\eta^{2}+\frac{12}{t}\int_{\Omega}F^{t+1}|\nabla\eta|^{2}.$
Now, choose $\eta_{1}\in C_{0}^{\infty}(\Omega)$ such that
$\begin{cases}0\leq\eta_{1}\leq 1,\quad\eta_{1}\equiv 1\penalty 10000\ in\penalty 10000\ B\left(o,\frac{3}{4}R\right);\\[8.53581pt] |\nabla\eta_{1}|\leq\frac{C(n)}{R}.\end{cases}$
and let $\eta=\eta_{1}^{t+2}$. Direct calculation shows that
$|\nabla\eta|^{2}\leq\frac{C^{2}(n)}{R^{2}}(t+2)^{2}\eta^{\frac{2(t+1)}{t+2}}.$
By Young inequality, we obtain
(4.16) $\begin{split}\frac{12}{t}\int_{\Omega}F^{t+1}|\nabla\eta|^{2}\leq&\frac{C^{2}(n)}{R^{2}}\frac{12(t+2)^{2}}{t}\int_{\Omega}F^{t+1}\eta^{\frac{2(t+1)}{t+2}}\\ \leq&\frac{C_{0}\varepsilon^{\theta}}{2}\int_{\Omega}F^{t+2}\eta^{2}+C_{1}(n,q,t)\varepsilon^{-\theta(t+1)}\frac{V}{R^{2(t+2)}},\end{split}$
where $C_{1}(n,q,t)$ is positive and depends only on $n$, $q$ and $t$.
Set
$\widetilde{\Omega}=\left\{x\in\Omega:F\geq\frac{4M_{0}}{C_{0}}\varepsilon^{q-\theta-1}\right\},$
then we have
(4.17) $\begin{split}2M_{0}\varepsilon^{q-1}\int_{\Omega}F^{t+1}\eta^{2}=&2M_{0}\varepsilon^{q-1}\int_{\widetilde{\Omega}}F^{t+1}\eta^{2}+2M_{0}\varepsilon^{q-1}\int_{\Omega\setminus\widetilde{\Omega}}F^{t+1}\eta^{2}\\ \leq&\frac{C_{0}\varepsilon^{\theta}}{2}\int_{\widetilde{\Omega}}F^{t+2}\eta^{2}+2M_{0}\varepsilon^{q-1}\int_{\Omega\setminus\widetilde{\Omega}}\left(\frac{4M_{0}}{C_{0}}\varepsilon^{q-\theta-1}\right)^{t+1}\\ \leq&\frac{C_{0}\varepsilon^{\theta}}{2}\int_{\Omega}F^{t+2}\eta^{2}+C_{2}(n,q,t)\varepsilon^{(q-1)+(q-\theta-1)(t+1)}V,\end{split}$
where $C_{2}(n,q,t)$ is a positive and depends only on $n$, $q$ and $t$.
Substituting (4.16) and (4.17) into (4.15) yields
(4.18) $\begin{split}H(t)\left\|F^{t+1}\eta^{2}\right\|_{L^{\chi}(\Omega)}\leq C_{1}(n,q,t)\varepsilon^{-\theta(t+1)}\frac{V}{R^{2(t+2)}}+C_{2}(n,q,t)\varepsilon^{(q-1)+(q-\theta-1)(t+1)}V.\end{split}$
From the definition of $F$, we conclude that
$H(t)\left\{\int_{B\left(o,\frac{3}{4}R\right)}\left[(v+\varepsilon)^{-\theta}\left(\frac{|\nabla\omega|^{2}}{\omega^{2}}+dv^{q-1}\right)\right]^{(t+1)\chi}\right\}^{\frac{1}{\chi}}$
$\leq$ $C_{1}(n,q,t)\varepsilon^{-\theta(t+1)}\frac{V}{R^{2(t+2)}}+C_{2}(n,q,t)\varepsilon^{(q-1)+(q-\theta-1)(t+1)}V.$
Hence,
$H(t)\left\{\int_{B\left(o,\frac{3}{4}R\right)}\left[(v+\varepsilon)^{-\theta}v^{q-1}\right]^{(t+1)\chi}\right\}^{\frac{1}{\chi}}\leq C_{3}(n,p,t)\left(\frac{1}{R^{2(t+2)}\varepsilon^{\theta(t+1)}}+\varepsilon^{(q-1)+(q-\theta-1)(t+1)}\right)V,$
where $C_{3}(n,q,t)$ is a positive number and depends only on $n$, $q$ and $t$. We finish the proof of Lemma 4.4. ∎
Now, we are ready to prove Theorem 1.5 in the case $n\geq 4$ and $\frac{n+3}{n-1}\leq q<\frac{n+2}{n-2}$.
Proof of Theorem 1.5 for the case $n\geq 4$ &\& $\frac{n+3}{n-1}\leq q<\frac{n+2}{n-2}$:.
Since $\theta\in\left(0,\,\frac{2}{n-2}\right)$, we have
$q-\theta-1>0.$
Next, we choose a large $t$ such that, as $\chi\in(1,\frac{n}{n-2})$,
(4.19) $t+3-\frac{2\chi}{\chi-1}>0\quad\mbox{and}\quad\frac{1}{\theta}\left[(q-1)+(q-\theta-1)(t+1)\right]-\frac{2\chi}{\chi-1}>0,$
or as $\chi=\frac{n}{n-2}$,
(4.20) $t+3-\beta^{*}>0\quad\mbox{and}\quad\frac{1}{\theta}\left[(q-1)+(q-\theta-1)(t+1)\right]-\beta^{*}>0,$
and the condition (4.10) holds. Once these have been done, we see that there exists a positive constant $C(n,q,\beta^{*})$ depending on $n$, $q$ and $\beta^{*}$ such that, if
$\|\mathrm{Ric}_{-}\|_{L^{\frac{\chi}{\chi-1}}}\leq C(n,q,\beta^{*})\mathbb{S}_{\chi}(M),$
then Lemma 4.4 holds.
By the fact that $\varepsilon\in(0,1)$, we infer from Lemma 4.4 that
$H(t)\left\{\int_{B\left(o,\frac{3}{4}R\right)}\left[(v+1)^{-\theta}v^{q-1}\right]^{(t+1)\chi}\right\}^{\frac{1}{\chi}}\leq C_{3}\left(\frac{1}{R^{2(t+2)}\varepsilon^{\theta(t+1)}}+\varepsilon^{(q-1)+(q-\theta-1)(t+1)}\right)V,$
where $\varepsilon$, $\theta$, $t$ and $H(t)$ are defined in Lemma 4.4.
Let
$\varepsilon=R^{-\frac{1}{\theta}}\quad(R\geq 1),$
then we have
$H(t)\left\{\int_{B\left(o,\frac{3}{4}R\right)}\left[(v+1)^{-\theta}v^{q-1}\right]^{(t+1)\chi}\right\}^{\frac{1}{\chi}}\leq C_{3}\left(\frac{1}{R^{t+3}}+\frac{1}{R^{\frac{1}{\theta}\left[(q-1)+(q-\theta-1)(t+1)\right]}}\right)V.$
Since
$\mathrm{vol}\left(B(o,R)\right)=O(R^{\frac{2\chi}{\chi-1}})\quad\mbox{or}\quad\mathrm{vol}\left(B(o,R)\right)=O(R^{\beta^{*}}),$
by setting $\beta^{*}_{\chi}=\frac{2\chi}{\chi-1}$ as $1<\chi<\frac{n}{n-2}$ and $\beta^{*}_{\chi}=\beta^{*}$ as $\chi=\frac{n}{n-2}$, we obtain
(4.21) $H(t)\left\{\int_{B\left(o,\frac{3}{4}R\right)}\left[(v+1)^{-\theta}v^{q-1}\right]^{(t+1)\chi}\right\}^{\frac{1}{\chi}}\leq C_{4}\left(\frac{1}{R^{t+3-\beta^{*}_{\chi}}}+\frac{1}{R^{\frac{1}{\theta}\left[(q-1)+(q-\theta-1)(t+1)\right]-\beta^{*}_{\chi}}}\right).$
Combining (4.20) and (4.21) together and letting $R\longrightarrow+\infty$, we arrive at
$\left\{\int_{M}\left[(v+1)^{-\theta}v^{q-1}\right]^{(t+1)\chi}\right\}^{\frac{1}{\chi}}=0.$
Therefore, $v\equiv 0$. This contradicts to the fact $v$ is a positive solution. We complete the proof. ∎
4.B. Case 2: $n\in\{2,3\}\quad\&\quad q\in\left[\frac{n+3}{n-1},\penalty 10000\ \frac{n+2}{(n-2)_{+}}\right)$
4.B.1. Second auxiliary function $G$ and estimation of the leading coefficients.
For $\theta\neq 0$, let $\omega=(v+\varepsilon)^{-\theta}$, then we have,
(4.22) $\Delta\omega=\left(1+\frac{1}{\theta}\right)\frac{|\nabla\omega|^{2}}{\omega}+\theta\omega\frac{v^{q}}{v+\varepsilon}.$
For undetermined real numbers $\varepsilon>0$ and $d>0$, define the second type auxiliary function:
(4.23) $G=(v+\varepsilon)^{-\theta}\left(\frac{|\nabla\omega|^{2}}{\omega^{2}}+dv^{q-1}\right).$
Lemma 4.5 (Lemma 2.2 and Lemma 6.5 in [47]).
There holds:
$\begin{split}\omega^{-1}\Delta G=&2\omega^{-2}\left|\nabla^{2}\omega-\frac{\Delta\omega}{n}g\right|^{2}+2\omega^{-2}\mathrm{Ric}(\nabla\omega,\nabla\omega)+\frac{2}{\theta}\omega^{-1}\langle\nabla G,\nabla\ln\omega\rangle\\ &+U\frac{|\nabla\omega|^{4}}{\omega^{4}}+V\frac{|\nabla\omega|^{2}}{\omega^{2}}v^{q-1}+Wv^{2(q-1)},\end{split}$
where
$U=$ $\left[\frac{2}{n}\left(1+\frac{1}{\theta}\right)-1\right]\left(1+\frac{1}{\theta}\right),$
$V=$ $\left[\frac{4}{n}(1+\theta)+2+\theta\right]\frac{v}{v+\varepsilon}-2q+\frac{2d}{\theta}\left(\frac{q-1}{\theta}\frac{v+\varepsilon}{v}-1\right)$
$+d\left[\frac{(q-1)(q-2)}{\theta^{2}}\frac{(v+\varepsilon)^{2}}{v^{2}}+1+\frac{1}{\theta}-\frac{2(q-1)}{\theta}\frac{v+\varepsilon}{v}\right],$
$W=$ $\frac{2\theta^{2}}{n}\frac{v^{2}}{(v+\varepsilon)^{2}}+d\left(\frac{\theta v}{v+\varepsilon}+1-q\right).$
Moreover, for $n\in\{2,3\}$ and $q\in\left[\frac{n+3}{n-1},\penalty 10000\ \frac{n+2}{(n-2)_{+}}\right)$, there exist constants $\theta=\theta(n,q)\in\left(0,\,q-1\right)$ if $n=2$ or $\theta=\theta(n,q)\in\left(0,\,\min\left\{2,q-1\right\}\right)$ if $n=3$, $d=d(n,q)>0$, $L=L(n,q)>0$ and $\widetilde{M}=\widetilde{M}(n,q)>0$ such that for any $\varepsilon>0$,
$U$ $\geq U_{0}>0,$
$V$ $\geq V_{0}-\widetilde{M}\overline{\chi}_{\left\{x\in B(o,R):\,v(x)<L\varepsilon\right\}},$
$W$ $\geq W_{0}-\widetilde{M}\overline{\chi}_{\left\{x\in B(o,R):\,v(x)<L\varepsilon\right\}},$
where $U_{0}$, $V_{0}$ and $W_{0}$ are positive constants depending only on $n$ and $q$.
Proof of Theorem 1.5 for the case $n\in\{2,3\}$ &\& $\frac{n+3}{n-1}\leq q<\frac{n+2}{(n-2)_{+}}$:.
In the case $n\in\{2,3\}$ and $q$ satisfies
$\frac{n+3}{n-1}\leq q<\frac{n+2}{(n-2)_{+}},$
the proof of Theorem 1.5 goes almost the same as that in the case $n\geq 4$ and $\frac{n+3}{n-1}\leq q<\frac{n+2}{n-2}$. Now, we sketch the proof here.
Following the lines of proof of Lemma 4.2, we obtain there exist $C_{0}=C_{0}(n,q)>0$ and $M_{0}=M_{0}(n,q)>0$, such that for any $\varepsilon>0$, there holds
$\Delta G\geq-2\mathrm{Ric}_{-}G-\frac{2}{\theta}|\nabla G||\nabla\ln\omega|+C_{0}(v+\varepsilon)^{\theta}G^{2}-M_{0}\varepsilon^{q-1}G.$
Then, following the lines of proof of Lemma 4.3, we obtain that, for
$t\in\left(\max\left\{\frac{8}{C_{0}\theta^{2}},1\right\},\penalty 10000\ +\infty\right),$
the following holds
$\begin{split}&C_{0}\varepsilon^{\theta}\int_{\Omega}G^{t+2}\eta^{2}+\frac{\mathbb{S}_{\chi}(M)}{2t}\left\|G^{t+1}\eta^{2}\right\|_{L^{\chi}(\Omega)}\\ \leq&4\int_{\Omega}\mathrm{Ric}_{-}G^{t+1}\eta^{2}+2M_{0}\varepsilon^{q-1}\int_{\Omega}G^{t+1}\eta^{2}+\frac{12}{t}\int_{\Omega}G^{t+1}|\nabla\eta|^{2},\end{split}$
where $\eta\geq 0$ and $\eta\in C_{0}^{\infty}(\Omega)$.
Once this has been done, it follows from the proof of Lemma 4.4 that, if
$H(t)=\frac{\mathbb{S}_{\chi}(M)}{2t}-4\left\|\mathrm{Ric_{-}}\right\|_{L^{\frac{\chi}{\chi-1}}}>0,$
then there exists $C_{3}=C_{3}(n,q,t)>0$ such that
$H(t)\left\{\int_{B\left(o,\frac{3}{4}R\right)}\left[(v+1)^{-\theta}v^{q-1}\right]^{(t+1)\chi}\right\}^{\frac{1}{\chi}}\leq C_{3}\left(\frac{1}{R^{2(t+2)}\varepsilon^{\theta(t+1)}}+\varepsilon^{(q-1)+(q-\theta-1)(t+1)}\right)V,$
where $V$ denotes the volume of geodesic ball $B(o,R)$ and $\varepsilon\in(0,1)$ is any positive number.
Finally, following the lines of proof of Theorem 1.5 for the case $n\geq 4$ and $\frac{n+3}{n-1}\leq q<\frac{n+2}{n-2}$ we can finish the proof.∎
Proof of Theorem 1.5:.
Combining Theorem 1.4, and the conclusions in this section, we finish the proof of Theorem 1.5. ∎
5. Geometric applications
Now we want to make use of some conclusions obtain in the previous sections to study the topological properties of a manifold which enjoy a Sobolev inequality. As mentioned in the previous, many manifolds, for instances Riemannian manifold with $\mathrm{Ric}\geq 0$ and $0<\mathrm{AVR}_{g}\leq 1$, minimal hypersurfaces [49] and Ricci shrinker [45, 66], support Sobolev inequalities. In particular, Brendle [8] obtained the sharp Sobolev inequality on complete noncompact Riemannian manifolds with nonnegative Ricci curvature using the so-called ABP method, addressing an open question posed by Cordero-Erausquin, Nazaret, and Villani [24] for such manifold. Balogh and Kristály [4] provided an alternative proof of Brendle’s rigidity result and confirmed that the Sobolev constant is sharp. Concretely, their results can be stated as follows:
Let $\left(M^{n},g\right)$ be a noncompact, complete $n$-dimensional Riemannian manifold with $\mathrm{Ric}\geq 0$ and $0<\mathrm{AVR}_{g}\leq 1$, where $\mathrm{Ric}$ denote the Ricci curvature of $(M,g)$. Then for all $v\in C_{0}^{\infty}\left(M^{n}\right)$, there holds
$\left\lVert v\right\rVert_{L^{2n/(n-2)}\left(M^{n}\right)}\leq\mathcal{S}\left(\mathbb{R}^{n}\right)\mathrm{AVR}_{g}^{-1/n}\left\lVert\nabla v\right\rVert_{L^{2}\left(M^{n}\right)}.$
Furthermore, the constant $\mathcal{S}\left(\mathbb{R}^{n}\right)\mathrm{AVR}_{g}^{-1/n}$ is sharp.
Here, $\mathrm{AVR}_{g}$ is the asymptotic volume ratio of $\left(M^{n},g\right)$, defined as
$\mathrm{AVR}_{g}=\lim_{R\to\infty}\dfrac{\mathrm{Vol}(B_{R}(o))}{\omega_{n}R^{n}},$
and $\omega_{n}$ is the volume of the unit ball in $\mathbb{R}^{n}$ and $\mathcal{S}\left(\mathbb{R}^{n}\right)$ is the best Sobolev constant in $\mathbb{R}^{n}$.
Using harmonic function theory to study geometric and topological properties of manifolds has a long history, for more details we refer to [42, 43, 67] and references therein. Here, firstly we need to recall the definition of end in a complete noncompact manifold:
Definition 5.1.
Let $(M,g)$ be a complete noncompact Riemannian manifold. We say that $M$ has $k$ ends $E_{1},\cdots,E_{k}$, if there exists a bounded open set $U$ so that
$M\setminus\overline{U}=\cup_{i=1}^{k}E_{i}$
with each $E_{i}$ being a noncompact connected component of $M\setminus\overline{U}$. Usually, we also call $E_{1},\cdots,E_{k}$ as the ends corresponding to $U$.
Denote usually the linear space spanned by bounded harmonic functions on $(M,g)$ by $H^{\infty}(M)$. Now let’s recall some previous conclusions. The first named author of this paper has ever shown the following:
Theorem 5.2 ([67], Theorem 3.3).
Let $(M,g)$ be a complete noncompact Riemannian manifold with $\dim(M)\geq 3$, Sobolev constant $\mathbb{S}_{\frac{n}{n-2}}(M)>0$ and $\mathrm{Ric}(M)\geq 0$ outside some compact subset. Then $M$ has only finitely many ends $E_{1},E_{2},\cdots,E_{k}$ and $\dim H^{\infty}(M)=k$.
In fact, this theorem and the Cheng-Yau’s gradient estimate of positive harmonic functions (see [17]) imply that a complete noncompact Riemannian manifold, which satisfies $n=\dim(M)\geq 3$, Sobolev constant $\mathbb{S}_{\frac{n}{n-2}}(M)>0$ and $\mathrm{Ric}(M)\geq 0$, has only an end.
The philosophy of the proof of the above theorem is: if $(M,g)$ has at least two ends and the Sobolev constant of $(M,g)$ is positive, then there exists a nonconstant, bounded, and positive harmonic function on $(M,g)$. Later on, this conclusion was also derived in [12].
As an extension of Theorem 5.2, we can take almost the same arguments as in [30] and [67] to conclude the following theorem, i.e., Theorem 1.10.
Theorem 5.3.
Let $(M,g)$ be a complete noncompact Riemannian manifold with Sobolev constant $\mathbb{S}_{\chi}(M)>0$. If $(M,g)$ has at least $k$ ends, then $\dim(H^{\infty}(M))\geq k$.
Proof.
Actually, by taking almost the same arguments as in [30] and [67], we can show the theorem. For the sake of completeness, here we give the routine to construct bounded positive harmonic functions on such a manifold $(M,g)$ which has at least $k$ ends and the Sobolev constant $\mathbb{S}_{\chi}(M)$ of $(M,g)$ is of a positive lower bound.
We let $\Omega_{i}$ ($i=1,2,\cdots$) be an exhaustion of $(M,g)$, i.e., each $\Omega_{i}$ is an open domain contained in $M$ and $\overline{\Omega}_{i}$ is compact, $\overline{\Omega}_{i}\subset\Omega_{i+1}$ for every $i\geq 1$, and $\cup_{i=1}^{\infty}\Omega_{i}=M$. For $i\geq i_{0}$ and $i_{0}$ sufficiently large, let
$M\setminus\Omega_{i}=\cup^{s}_{j=1}E^{(i)}_{j},\quad s\geq k$
be the disjoint union of connected components. Pick $k$ ends of $\{E^{(i)}_{j}:1\leq j\leq s\}$ and denote the $k$ ends of $(M,g)$ by $E_{1}^{(i_{0})},\cdots,E_{k-1}^{(i_{0})}$ and $E_{k}^{(i_{0})}$, and for the sake of convenience, by $E_{1},\cdots,E_{k}$ without confusions.
Moreover, denote $E_{j}^{i}=E_{j}\cap\Omega_{i}$. Now we consider the following Dirichlet problems of harmonic functions:
$\Delta u_{j}^{i}=0,\quad u=0\,\,\mbox{on}\,\,\partial\Omega_{i}\setminus\partial E_{j}^{i}\quad\mbox{and}\quad u_{j}^{i}=1\,\,\mbox{on}\,\,\partial E_{j}^{i},$
where $j=1,\cdots,k$. Thus we can obtain $k$ sequences of harmonic functions, denoted as $\{u_{j}^{i}\}$ ($j=1,2,\cdots,k$).
As done in [30], we let $\varphi_{j}(x)$ be a real smooth function on $(M,g)$ such that
$\varphi_{j}(x)\equiv 1\quad\mbox{on}\,\,E_{j}^{(i_{0})}\quad\mbox{and}\quad\varphi_{j}(x)\equiv 0\quad\mbox{in}\,\,M\setminus(\Omega_{i_{0}}\cup E_{j}^{(i_{0})});$
and then define $\rho_{j,i}$ on $\Omega_{i}$ by
$\rho_{j,i}=|u^{i}_{j}-\varphi_{j}|\quad\mbox{for}\,\,i\geq i_{0}.$
It is easy to see that $|\Delta\varphi_{j}|\in L^{p}$ for any $p>0$ and in the sense of distribution (see [30])
$\Delta\rho_{j,i}\geq-|\Delta\varphi_{j}|.$
We choose a test function $\rho_{j,i}^{\alpha-1}$ ($\alpha>1$) to multiply the above equation and integrate on $\Omega_{i}$
$\int_{\Omega_{i}}\rho_{j,i}^{\alpha-1}\Delta\rho_{j,i}\geq-\int_{\Omega_{i}}\rho_{j,i}^{\alpha-1}|\Delta\varphi_{j}|.$
Then we have
$\frac{4(\alpha-1)}{\alpha^{2}}\int_{\Omega_{i}}\left|\nabla\rho_{j,i}^{\frac{\alpha}{2}}\right|^{2}\leq\int_{\Omega_{i}}\rho_{j,i}^{\alpha-1}|\Delta\varphi_{j}|.$
Since $(M,g)$ enjoys a Sobolev inequality as follows
$\mathbb{S}_{\chi}(M)\left(\int_{M}f^{2\chi}dv\right)^{\frac{1}{\chi}}\leq\int_{M}|\nabla f|^{2}dv,$
we obtain
$\left(\int_{\Omega_{i}}\rho_{j,i}^{\alpha\chi}\right)^{\frac{1}{\chi}}\leq\frac{\alpha^{2}}{4(\alpha-1)}\mathbb{S}_{\chi}(M)^{-1}\int_{\Omega_{i}}\rho_{j,i}^{\alpha-1}|\Delta\varphi_{j}|.$
Applying the Hölder inequality we can see
$\left(\int_{\Omega_{i}}\rho_{j,i}^{\alpha\chi}\right)^{\frac{1}{\chi}}\leq C\left(\int_{\Omega_{i}}\rho_{j,i}^{\alpha\chi}\right)^{\frac{\alpha-1}{\alpha\chi}}\left(\int_{\Omega_{i}}|\Delta\varphi_{j}|^{\frac{\alpha\chi}{\alpha\chi-\alpha+1}}\right)^{\frac{\alpha\chi-\alpha+1}{\alpha\chi}}.$
It follows that
(5.1) $\left(\int_{\Omega_{i}}\rho_{j,i}^{\alpha\chi}\right)^{\frac{1}{\alpha\chi}}\leq C\left(\int_{\Omega_{i}}|\Delta\varphi_{j}|^{\frac{\alpha\chi}{\alpha\chi-\alpha+1}}\right)^{\frac{\alpha\chi-\alpha+1}{\alpha\chi}}.$
By the same arguments as in [67] and Corollary 4.3 of [30] we can see that there exists $k$ harmonic functions $u_{1},\cdots,u_{k-1}$ and $u_{k}$ on $(M,g)$ such that by neglecting some subsequences $u_{1}^{i},\cdots,u_{k-1}^{i}$ and $u_{k}^{i}$ converge respectively in the sense of $C^{k}$ ($k\geq 2$) to $u_{1},\cdots,u_{k-1}$ and $u_{k}$ on any compact subset of $(M,g)$, where $j=1,\cdots,k$, i.e.,
$\lim_{i\to\infty}\|u^{i}_{j}-u_{j}\|_{C^{2}(\mathcal{C})}=0\quad\quad j=1,\cdots,k$
where $\mathcal{C}\in M$ is any compact subset. In fact, we can also use directly Cheng-Yau’s local gradient estimate [17] to show the convergence of these sequences $\{u^{i}_{j}\}$ ($j=1,\cdots,k;\,\,i=1,2,\cdots$).
Now, we claim that $u_{1},\cdots,u_{k-1}$ and $u_{k}$ are linearly independent in the linear space spanned by bounded harmonic functions on $(M,g)$, i.e., $\dim(H^{\infty}(M))\geq k$.
In order to prove the above assertion, firstly we note that by maximum principle there hold true $0<u_{j}(x)<1$ for any $x\in M$, where $j=1,2,\cdots,k$. Then, we let $i\to\infty$ in (5.1)and make use of a diagonal technique to obtain
(5.2) $\left(\int_{E_{j}}(1-u_{j})^{\alpha\chi}\right)^{\frac{1}{\alpha\chi}}<\infty\quad\quad\mbox{and}\quad\quad\left(\int_{E_{l}}u_{j}^{\alpha\chi}\right)^{\frac{1}{\alpha\chi}}<\infty\quad\mbox{for}\,\,l\neq j,$
since we have
$\int_{M}|\Delta\varphi_{j}|^{\frac{\alpha\chi}{\alpha\chi-\alpha+1}}<\infty.$
Therefore, it is easy to verify from the above integral estimates that $u_{1},\cdots,u_{k-1}$ and $u_{k}$ are not constant functions, since ${\rm{Vol}}(E_{j})=\infty$ for each $j=1,\cdots,k$.
The remaining is to show that $u_{1},\cdots,u_{k-1}$ and $u_{k}$ are linearly independent. Since for any geodesic ball $B_{r}\subset M$ we have the volume estimate
$\mathrm{Vol}\left(B_{r}\right)\geq C\left(\chi,\mathbb{S}_{\chi}(M)\right)r^{\frac{2\chi}{\chi-1}},$
this tells us that $\mathrm{Vol}(B_{r}(x))$ is of a uniform lower bound. Hence, noting $\chi>1$ and $\alpha>1$ we have
$\frac{1}{\mathrm{Vol}(B_{r}(x))}\int_{B_{r}(x)}u_{j}\leq\left(\frac{1}{\mathrm{Vol}(B_{r}(x))}\int_{B_{r}(x)}u_{j}^{\alpha\chi}\right)^{\frac{1}{\alpha\chi}}.$
It follows from (5.2) that for $B_{r}(x)\subset E_{i}$, where $r>0$ is a fixed number, and $j\neq i$ there hold true
$\lim_{x\in E_{i}\to\infty}\mathrm{Vol}(B_{r}(x))^{-1}\int_{B_{r}(x)}u_{j}=\lim_{x\in E_{i}\to\infty}\mathrm{Vol}(B_{r}(x))^{-\frac{1}{\alpha\chi}}\left(\int_{B_{r}(x)}u_{j}^{\alpha\chi}\right)^{\frac{1}{\alpha\chi}}=0$
and
$\lim_{x\in E_{j}\to\infty}\mathrm{Vol}(B_{r}(x))^{-1}\int_{B_{r}(x)}(1-u_{j})=\lim_{x\in E_{j}\to\infty}\mathrm{Vol}(B_{r}(x))^{-\frac{1}{\alpha\chi}}\left(\int_{B_{r}(x)}(1-u_{j})^{\alpha\chi}\right)^{\frac{1}{\alpha\chi}}=0.$
If there holds
$\sum_{j=1}^{k}a_{j}u_{j}(x)\equiv 0\quad\mbox{on}\,\,M,$
where $a_{j}$ ($j=1,2,\cdots,k$) are some real numbers, then we have
$\lim_{x\in E_{1}\to\infty}\mathrm{Vol}(B_{r}(x))^{-1}\int_{B_{r}(x)}\sum_{j=1}^{k}a_{j}u_{j}=0,$
which implies
$\lim_{x\in E_{1}\to\infty}\mathrm{Vol}(B_{r}(x))^{-1}\int_{B_{r}(x)}a_{1}u_{1}=0.$
We conclude that $a_{1}=0$. Otherwise, we have
$\lim_{x\in E_{1}\to\infty}\mathrm{Vol}(B_{r}(x))^{-1}\int_{B_{r}(x)}u_{1}=0.$
On the other hand, by the integral estimates (5.2) we have
$\lim_{x\in E_{1}\to\infty}\mathrm{Vol}(B_{r}(x))^{-1}\int_{B_{r}(x)}(1-u_{1})=0.$
Hence, it follows
$\lim_{x\in E_{1}\to\infty}\mathrm{Vol}(B_{r}(x))^{-1}\int_{B_{r}(x)}(1-u_{1})+\lim_{x\in E_{1}\to\infty}\mathrm{Vol}(B_{r}(x))^{-1}\int_{B_{r}(x)}u_{1}=0.$
However, we have that for any $B_{r}(x)\subset E_{1}$ there holds
$1=\mathrm{Vol}(B_{r}(x))^{-1}\int_{B_{r}(x)}1=\mathrm{Vol}(B_{r}(x))^{-1}\int_{B_{r}(x)}\{(1-u_{1})+u_{1}\}$
which implies
$\lim_{x\in E_{1}\to\infty}\mathrm{Vol}(B_{r}(x))^{-1}\int_{B_{r}(x)}(1-u_{1})+\lim_{x\in E_{1}\to\infty}\mathrm{Vol}(B_{r}(x))^{-1}\int_{B_{r}(x)}u_{1}=1.$
This is a contradiction. Similarly, we can also show that $a_{2}=0,\cdots,a_{k-1}=0$ and $a_{k}=0$. This indicates that $u_{1},\cdots,u_{k-1}$ and $u_{k}$ are linearly independent, hence we have shown
$\dim(H^{\infty}(M))\geq k.$
Thus, we finish the proof. ∎
Remark 5.4.
In fact, Theorem 5.3 can be extended to the following case $(M,g)$ enjoys a more general Sobolev inequality:
$\mathbb{S}_{\chi}(M)\left(\int_{M}f^{2\chi}dv\right)^{\frac{1}{\chi}}\leq\int_{M}(|\nabla f|^{2}+Rf^{2})dv,\quad\mbox{for any}\,\,f\in C_{0}^{\infty}(M),$
which appears in the geometry of submanifolds and Ricci solitons (see [45, 49]). Here $R(x)\geq 0$ is usually a smooth function related to the curvature. Obviously, it is easy to verify the inequality reduces to
$\mathbb{C}_{\chi}(M)\left(\int_{M}f^{2\chi}dv\right)^{\frac{1}{\chi}}\leq\int_{M}|\nabla f|^{2}dv,\quad\mbox{for any}\,\,f\in C_{0}^{\infty}(M),$
provided $\|R\|_{L^{\frac{\chi}{\chi-1}}}<\mathbb{S}_{\chi}(M)$. Particularly, such inequality with $\chi=\frac{n}{n-2}$ holds true in a complete minimal hypersurface in an Euclidean space (see [8, 49]).
Proof of Theorem 1.11: We prove this theorem by contradiction. If $(M,g)$ has at least two ends, then, by the assumption the $\frac{n}{n-2}$-type Sobolev constant $\mathbb{S}_{\frac{n}{n-2}}(M)$ is positive and Theorem 1.10 we conclude that there exists a nonconstant, bounded, and positive harmonic function on $(M,g)$ (see also Theorem B in [67] and Corollary 4.3 in [30] for more details). However, Corollary 1.7 tells us that there exists a positive constant $C(n,\beta^{*})$ depending on $n$ and $\beta^{*}$ such that, if
$\|\mathrm{Ric}_{-}\|_{L^{\frac{n}{2}}}\leq C(n,\beta^{*})\mathbb{S}_{\frac{n}{n-2}}(M),$
then there is no nonconstant, positive harmonic function on $(M,g)$. We obtain a contradiction. We complete the proof.
Theorem 5.5.
Let $(M^{n},g)$ be a complete noncompact Riemannian $n$-dimensional manifold with Sobolev constant $\mathbb{S}_{\chi}(M)>0$ where $1<\chi\leq\frac{n}{n-2}$. If ${\rm{Vol}}(B_{r}(x_{0}))=O\left(r^{\frac{2\chi}{\chi-1}}\right)$ for some geodesic ball $B_{r}(x_{0})\subset M$, then $M$ has only finitely many ends.
In particular, suppose $(M^{n},g)$ supports a $\chi$-type Sobolev inequality for some $\chi\in(1,n/(n-2))$, and assume in addition that $\|\mathrm{Ric}_{-}\|_{L^{\frac{\chi}{\chi-1}}}$ is bounded. Then $M$ has only finitely many ends.
Proof.
The basic idea of proof is contained in the arguments on Lemma 3.2 of [67]. For the sake of completeness, here we give the details. First, for any $r>0$ large enough we pick a maximal family of points $x_{i}\in\partial B_{r}(x_{0})$ such that $B_{\frac{r}{4}}(x_{i})$ are disjoint. Then we have
$\partial B_{r}(x_{0})\subset\cup_{i=1}^{q}\partial B_{\frac{r}{2}}(x_{i}),$
where $q=\sharp\{x_{j}\}$. In fact, if there exists some $\overline{x}\in\partial B_{r}(x_{0})\setminus\cup_{i=1}^{q}B_{\frac{r}{2}}(x_{i})$, then $B_{\frac{r}{4}}(\overline{x})$ is disjoint with $B_{\frac{r}{4}}(x_{i})$ for $i=1,\cdots,q$. This contradicts the fact that $\{x_{i}\}$ is a maximal family.
Obviously we have the following inequality:
$\sum_{i=1}^{q}{\rm{Vol}}(B_{\frac{r}{4}}(x_{i}))\leq{\rm{Vol}}(B_{2r}(x_{0})).$
By the assumption of this theorem, we have
${\rm{Vol}}(B_{2r}(x_{0}))\leq C(n)r^{\frac{2\chi}{\chi-1}}.$
By Theorem 1.3 we have
${\rm{Vol}}(B_{\frac{r}{4}}(x_{i}))\geq C\left(\chi,\mathbb{S}_{\chi}(M)\right)r^{\frac{2\chi}{\chi-1}},$
Hence, combining the above inequalities, we obtain
$qC\left(\chi,\mathbb{S}_{\chi}(M)\right)\leq C(n).$
It follows that
$q<\left[\frac{C(n)}{C\left(\chi,\mathbb{S}_{\chi}(M)\right)}\right]+1\equiv\Lambda,$
where $C(n)/C\left(\chi,\mathbb{S}_{\chi}(M)\right)]$ is the integer part of $C(n)/C\left(\chi,\mathbb{S}_{\chi}(M)\right)$. This means that the counting number of the above maximal family denoted by $\sharp\{x_{j}\}$ is bounded by a constant depending only on $n$, $\chi$ and $\mathbb{S}_{\chi}(M)$.
We claim that $M$ has only finitely many ends denoted by $E_{1},\cdots,E_{k-1}$ and $E_{k}$ and $k\leq\Lambda$. Otherwise, we choose $r_{0}>0$ such that
$M\setminus B_{r_{0}}(x_{0})=\cup^{s}_{j=1}E^{(r_{0})}_{j},\quad s\geq\Lambda+1$
is the disjoint union of connected components.
Now choose points $x_{j}^{\prime}\in\partial B_{4r_{0}}(x_{0})\cap E^{(r_{0})}_{j}$ for $j=1,\dots,s$. It is then straightforward to see that the balls $B_{r_{0}}(x_{j}^{\prime})$ $(j=1,\dots,s)$ are pairwise disjoint. Indeed, if $B_{r_{0}}(x_{i}^{\prime})\cap B_{r_{0}}(x_{j}^{\prime})\neq\varnothing$ for some $i\neq j$, then $d(x_{i}^{\prime},x_{j}^{\prime})<2r_{0}$. On the other hand, since $x_{i}^{\prime}$ and $x_{j}^{\prime}$ lie in different connected components, any minimizing geodesic joining them must intersect $\partial B_{r_{0}}(x_{0})$. Hence $d(x_{i}^{\prime},x_{j}^{\prime})\geq 6r_{0},$ a contradiction. Therefore the family $\{B_{r_{0}}(x_{j}^{\prime})\}_{j=1}^{s}$ is disjoint, which contradicts the fact that the cardinality of maximal disjoint family is at most $\Lambda$. We complete the proof. ∎
Acknowledgments: We would like to thank Professor Jun Yang for insightful discussions. Y. Wang is supported by National Natural Science Foundation of China (Grant No.12431003); G. Wei is supported by National Natural Science Foundation of China (Grants No.12101619 and 12141106).